### v2

# A phase transition for the two-dimensional random field Ising/FK-Ising model

Abstract
We study the total variation (TV) distance between the laws of the 2D Ising/FK-Ising model in a box of side-length $N$ with and without an i.i.d. Gaussian external field with variance $\epsilon^{2}$. Letting the external field strength $\epsilon=\epsilon(N)$ depend on the size of the box, we derive a phase transition for each model depending on the order of $\epsilon(N)$. For the random field Ising model, the critical order for $\epsilon$ is $N^{-1}$. For the random field FK-Ising model, the critical order depends on the temperature regime: for $T>T_{c}$, $T=T_{c}$ and $T\in(0,T_{c})$ the critical order for $\epsilon$ is, respectively, $N^{-\frac{1}{2}}$, $N^{-\frac{15}{16}}$ and $N^{-1}$. In each case, as $N\to\infty$ the TV distance under consideration converges to $1$ when $\epsilon$ is above the respective critical order and converges to 00 when below.
1 Introduction
The Ising model is one of the most significant models in statistical physics exhibiting the phenomenon of phase transition. The impact of minor perturbations on the phase transition phenomenon stands as an important question in disordered systems, and extensive study has been dedicated to exploring this topic within the context of the Ising model. Specifically, Imry-Ma [27] predicts that the long-range order—or, equivalently, a positive lower bound for the boundary influence (as defined subsequently)—exists for the random field Ising model (RFIM) with weak disorder when the temperature $T$ is small in dimensions three and higher, but not in two dimensions for all temperatures $T$. Regarding the two-dimensional behavior, the absence of long-range order was first proved in [3]. Some estimations on the decay rate of the boundary influence were first given in [2, 13], and it was finally shown in [1, 19] that the boundary influence decays exponentially, again at all temperatures. For dimensions $d\geq 3$, let $T_{c}(d)$ denote the critical temperature in $d$ dimensions of the usual Ising model, with no external field. The existence of long-range order was proved at low enough temperature and small enough disorder in [26, 7]. The existence of long-range order at low enough temperature was recently reproved in [20] and then extended to all temperatures below the critical temperature $T_{c}(d)$ in [16]. Additionally, a correlation inequality (from [17]) indicates that the long-range order does not exist for any temperatures at or above $T_{c}(d)$.
Despite the significant progress regarding the Imry-Ma predictions, some relevant properties of the RFIM are still not fully comprehended. One such property is the near-critical behavior of the two-dimensional RFIM. More specifically, examining what happens when the disorder strength $\epsilon$ varies with the box size $N$ at different rates. In this direction, [18], [20] and [15] study the correlation length—roughly speaking, when critical order that the boundary influence “perceive” the external field, see Section 2.6 for detailed definitions—of the two-dimensional RFIM for all $T\leq T_{c}.$ Furthermore, in [8, 9, 10], they study the Ising model with a uniform external field and show the phase transition of boundary influence with respect to the strength of the external field.
However, despite its significance, the boundary influence is merely a statistic of the random field Ising measure. This naturally leads us to ponder upon the following question: when does the measure itself “perceive” the external field? More formally, how does the total variation distance between the Ising measures with and without external field vary with $\epsilon$? The primary objective of this paper is to provide a definitive answer to this question.
1.1 The random field Ising model: a warm-up
Let $\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}$ and $\mu^{\xi,0}_{T,\Lambda_{N}}$ stand for the two-dimensional Ising measures (see Section 2.2 for detailed definitions) with and without random external field $\epsilon h$, where $h=\{h_{x}:x\in\mathbb{Z}^{d}\}$ is a family of i.i.d. standard normal variables with law $\mathbb{P}$. Our first result concerns the phase transition of the total variation (denoted by $\|\cdot\|_{\mathrm{TV}}$) distance between them.
Theorem 1.1.
For any temperature $T>0$, there exists a positive constant $c=c(T)$ such that for any box size $N$, external field strength $\epsilon>0$ and boundary condition $\xi$:
$\mathbb{P}\left(\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}>c^{-1}\sqrt{\epsilon N}\right)\leq c^{-1}\exp\left(-c\epsilon^{-\frac{1}{4}}N^{-\frac{1}{4}}\right).$ (1.1)
$\mathbb{P}\left(\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}<1-c^{-1}\sqrt{\epsilon^{-1}N^{-1}}\right)\leq c^{-1}\exp\left(-c\epsilon N\right).$ (1.2)
Writing $a(N)\ll b(N)$ if $\lim\limits_{N\rightarrow\infty}\frac{a(N)}{b(N)}=0$ and $a(N)\gg b(N)$ if $\lim\limits_{N\rightarrow\infty}\frac{a(N)}{b(N)}=\infty$, we have the following corollary immediately:
Corollary 1.2.
Fix any temperature $T>0,$ then for any boundary condition $\xi$:
  1. (i)
    if $\epsilon=\epsilon(N)\ll N^{-1}$, we have $\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}\rightarrow 0$ in probability as $N\to\infty$;
  2. (ii)
    if $\epsilon=\epsilon(N)\gg N^{-1}$, we have $\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}\rightarrow 1$ in probability as $N\to\infty$.
In a word, the total variation distance is negligible for $\epsilon\ll N^{-1},$ and significant for $\epsilon\gg N^{-1}.$ To heuristically understand this result, one may note that there are $N^{2}$ Gaussian variables in $\Lambda_{N}$ thus the fluctuation of the Radon-Nikodym derivative between the measures should be of order $\epsilon^{2}N^{2}.$
1.2 Main results
Thanks to the power of the Edward-Sokal coupling, the FK-Ising model frequently plays a central role in the study of the Ising model, even with the presence of a random external field. Our main result studies the behavior of the total variation distance between the two-dimensional FK-Ising model and its random field variation, where we call it the random field FK-Ising model (RFFKIM). In this setting, we observe a more interesting phenomenon, showing that the behavior differs according to the temperature regime. More precisely, let $T_{c}=T_{c}(2)$ denote the critical temperature of the two-dimensional Ising model. Let $\phi^{\gamma,\epsilon h}_{p,\Lambda_{N}}$ and $\phi^{\gamma,0}_{p,\Lambda_{N}}$ denote the FK-Ising measures with and without external field, where $\gamma$ stands for the boundary condition and $p=1-\exp(-\frac{2}{T})$ is the parameter for the FK-Ising model, see Section 2.3 for detailed definitions. Throughout this paper, we will keep the convention
$p=1-\exp\left(-\frac{2}{T}\right)$
thereby treating $p$ and $T$ as two interchangeable representations of a single parameter. We introduce the following notation for simplicity:
$\alpha(T)=\left\{\begin{aligned} &1~{}~{}~{}&0<T<T_{c}\,,\\ &\frac{15}{16}~{}~{}~{}&T=T_{c}\,,\\ &\frac{1}{2}~{}~{}~{}&T>T_{c}\,.\end{aligned}\right.$
Theorem 1.3.
For any temperature $T>0$ and corresponding parameter $p=1-\exp(-\frac{2}{T}),$ there exists a constant $c=c(T)>0$ such that for any boundary condition $\gamma$:
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}>c^{-1}\sqrt{\epsilon N^{\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right).$ (1.3)
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}<1-c^{-1}\sqrt{\epsilon^{-1}N^{-\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon N^{\alpha(T)}\right).$ (1.4)
Corollary 1.4.
Fix $T\in(0,+\infty)$. Then for any boundary condition $\gamma$, the followings hold:
  1. (i)
    if $\epsilon=\epsilon(N)\ll N^{-\alpha(T)}$, we have $\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}\rightarrow 0$ in probability as $N\to\infty$;
  2. (ii)
    if $\epsilon=\epsilon(N)\gg N^{-\alpha(T)}$, we have $\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}\rightarrow 1$ in probability as $N\to\infty$.
Remark 1.5.
We leave out the case $T=0$ in both Theorems 1.1 and 1.3 for the following reasons. For the former case, when $T=0$ the Ising measure is supported on the ground states which leads to a different phase transition point as $N^{-1}$ (see [18] for a related work). For the latter case, the definition of the FK-Ising model with external field does not directly extend to the $T=0$ case. Although in principle it is possible to give a definition using the Edward-Sokal coupling via the RFIM, the resulting object is not a natural extension of the FK-Ising model and is less interesting in our case.
1.3 Organization of the paper
In Section 2, we first present some notations and precise definitions of the RFIM and the RFFKIM, along with the fundamental properties of these models. Subsequently, in Section 3, we will prove Theorem 1.1 as a warm-up, since it reveals the core concepts of deriving upper and lower bounds of total variation distance through relatively straightforward calculations. Section 4 will be dedicated to proving the upper bound of the total variation distance between the FK-Ising model with and without external field (i.e. (1.3)). The corresponding lower bounds (i.e. (1.4)) will be postponed to Section 6, utilizing a coarse-graining framework to accumulate small differences on small boxes. Before that, we demonstrate that the differences in small boxes indeed exist due to the effect of the external field on boxes of an appropriate size in Section 5. In Appendix A, we prove an LDP result for the (discrete) FK-Ising model at critical temperature. While similar results exist for the continuous scaling limit [11, Theorem 3], the discrete case cannot be derived directly from it.
Throughout our proof, we have used several estimates on the FK-Ising model and Gaussian concentration inequalities, which may be considered intuitive or even elementary for experts in these fields; however, despite our best efforts, we could not locate precise references for these results. For the sake of completeness, we have included proofs of them in the Appendices B and C.
2 Preliminaries and notation
In this section, we introduce some basic definitions and notation.
2.1 Notation
We use $c,C,c_{i},C_{i}$ to represent positive constants whose actual values may vary from line to line. For a finite set $A$, we denote by $|A|$ the cardinality of $A$. We use $A^{c}$ to denote the complement of the set (or event) $A$. If $A$ is an event, we denote its indicator by $\mathbf{1}_{A}$. Throughout the paper, we use $\lfloor a\rfloor$ to denote the floor of a real number $a$, i.e., the largest integer less than or equal to $a$.
[…]
For two (discrete) probability measures $\mu$ and $\nu$ on the same measurable space $(\Xi,\mathcal{F})$, the total variation distance between them is defined as $\|\mu-\nu\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\omega\in\Xi}|\mu(\omega)-\nu(\omega)|.$
For simplicity, we denote $f(x)=\ln(\cosh(x))$ throughout the paper, where $\cosh(x)=\frac{e^{x}+e^{-x}}{2}$, as this function appears frequently.
2.2 The random field Ising model (RFIM)
For $\epsilon>0$, we define the random field Ising model (RFIM) Hamiltonian $H_{T,G}^{\xi,\epsilon h}$ on a finite subgraph $G=(V,E)$ of $\mathbb{Z}^{d}$ with the boundary condition ($\xi\in\{1,0,-1\}^{\mathbb{Z}^{d}})$, and with the external field $\epsilon h=\{\epsilon h_{v}:v\in\mathbb{Z}^{d}\}$ by
$H_{T,G}^{\xi,\epsilon h}(\sigma)=-\left(\sum_{{\begin{subarray}{c}u,v\in V\\ u\sim v\end{subarray}}}\sigma_{u}\sigma_{v}+\sum_{{\begin{subarray}{c}u\in V\\ v\notin V\\ u\sim v\end{subarray}}}\sigma_{u}\xi_{v}+\sum_{u\in V}\epsilon h_{u}\sigma_{u}\right)~{}~{}\mbox{for}~{}~{}\sigma\in\Sigma,$ (2.1)
where $\Sigma:=\{-1,1\}^{V}$. For $T>0$, we define $\mu_{T,G}^{\xi,\epsilon h}$ to be a Gibbs measure on $\Sigma$ at temperature $T$ by
$\mu_{T,G}^{\xi,\epsilon h}(\sigma)=\dfrac{1}{\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}}e^{-\frac{1}{T}H_{T,G}^{\xi,\epsilon h}(\sigma)}~{}~{}\mbox{for}~{}~{}\sigma\in\Sigma,$ (2.2)
where $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ is the partition function given by $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}=\sum_{\sigma\in\Sigma}e^{-\frac{1}{T}H_{T,G}^{\xi,\epsilon h}(\sigma)}.$ Note that $\mu_{T,G}^{\xi,\epsilon h}$ and $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ are random variables depending on the external field $\epsilon h$. We will denote by $\langle\cdot\rangle_{T,G}^{\xi,\epsilon h}$ the expectation taken with respect to $\mu_{T,G}^{\xi,\epsilon h}$. By taking $\epsilon=0$ we get the pure Ising model.
2.3 The random field FK-Ising model (RFFKIM)
In this subsection, we briefly review the FK-Ising model with external field, which was introduced in [23]. For a subgraph $G=(V,E)$ of $\mathbb{Z}^{d}$, we can consider an edge configuration $\omega\in\Xi=\{0,1\}^{E}$ where 00 indicates that the edge is closed and $1$ indicates the edge is open. Throughout this paper, a boundary condition $\gamma$ is an edge configuration on $G^{c}$, i.e., $\gamma\in\{0,1\}^{\mathrm{E}(\mathbb{Z}^{d})\setminus E},$ where $\mathrm{E}(\mathbb{Z}^{d})$ is the edge set of $\mathbb{Z}^{d}.$ We use $\mathrm{w}$ and $\mathrm{f}$ to denote the wired and free boundary conditions, that is, all the edges in $\gamma$ are open or closed, respectively.
For any $x\neq y\in V$, we say $x$ is connected to $y$ in $\omega$ under boundary condition $\gamma$ if there exists a sequence $x_{0}=x,x_{1},\cdots,x_{n}=y$ such that $\{x_{i},x_{i+1}\}\in\mathrm{E}(\mathbb{Z}^{d})$ and $\{x_{i},x_{i+1}\}$ is open for all $10\leq i\leq n-$ 1; this sequence is called an open path in $\omega$. So given any $\omega\in\Xi,$ the graph $G$ is divided into a disjoint union of connected components, we also call them open clusters. Let $\kappa(\omega^{\gamma})$ be the number of open clusters in $\omega$ under boundary condition $\gamma$. We use the notation $\mathfrak{C}(\omega^{\gamma})=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{\kappa(\omega^{\gamma})}\}$ to denote the collection of all clusters. For notation clarity, we drop the superscript $\gamma$ in $\omega^{\gamma}$ when $\gamma$ is clear in the context.
Then the FK-Ising model on $G$ with parameter $p\in(0,1)$ (and corresponding temperature $T\in(0,\infty)$ with $p=1-\exp(-\frac{2}{T})$) is a probability measure on $\Xi$ given by
$\phi^{\gamma,\epsilon h}_{p,G}(\omega)=\frac{1}{\mathcal{Z}_{\phi,p,G}^{\gamma,\epsilon h}}\prod_{e\in E}\hskip 5.0ptp^{\omega(e)}(1-p)^{1-\omega(e)}\prod_{j=1}^{\kappa(\omega^{\gamma})}\hskip 5.0pt2\cosh\left(\frac{\epsilon h_{\mathcal{C}_{j}}}{T}\right),$ (2.3)
where
$\mathcal{Z}_{\phi,p,G}^{\gamma,\epsilon h}=\sum_{\omega\in\Xi}\prod_{e\in E}\hskip 5.0ptp^{\omega(e)}(1-p)^{1-\omega(e)}\prod_{j=1}^{\kappa(\omega^{\gamma})}\hskip 5.0pt2\cosh\left(\frac{\epsilon h_{\mathcal{C}_{j}}}{T}\right)$
is the normalizing constant (partition function) of $\phi^{\gamma,\epsilon h}_{p,G}$ and $h_{\mathcal{C}_{j}}=\sum_{x\in\mathcal{C}_{j}}h_{x}.$ Similarly, we obtain the FK-Ising model without external field and the corresponding partition function by taking $\epsilon=0.$
For a subgraph $G=(V,E)$ of $\mathbb{Z}^{d}$, an edge configuration $\omega\in\Xi=\{0,1\}^{E}$ and a subset $E^{\prime}\subset E$, recall that $\omega_{|{E^{\prime}}}$ is the restriction of $\omega$ on $E^{\prime}$. In addition, if there exists $V^{\prime}\subset V$ such that $E^{\prime}=\{(x,y)\in E:x,y\in V^{\prime}\}$, then we also write $\omega_{|V^{\prime}}$ to denote the restriction of $\omega$ on $E^{\prime}$.
Throughout this paper, we focus on the case where $G$ is a subgraph of $\mathbb{Z}^{2}$, particularly when $G=\Lambda_{N}:=[-N,N]^{2}\cap\mathbb{Z}^{2}$, which represents a box of side length $2N$ centered at the origin. Note that the boundary condition $\gamma$ influences the measure $\phi^{\gamma,\epsilon h}_{p,\Lambda_{N}}$ only through the extra connections using edges outside $\Lambda_{N}$. Thus, we can also choose the boundary condition as a partition on the interior boundary $\partial_{int}\Lambda_{N}.$ The external field $\epsilon h$ will always be chosen as $\epsilon h=\{\epsilon h_{x}:x\in G\}$ where $h_{x}$’s are i.i.d. standard Gaussian variables which law will be denoted as $\mathbb{P}$ and $\epsilon$ stands for the strength of the external field. In this setup, the model will be called the two-dimensional random field FK-Ising model.
2.4 The Edward-Sokal coupling
The Edwards-Sokal coupling due to the paper [23], enables us to express the correlation functions of the Ising model in terms of the FK-Ising model. The Edward-Sokal coupling between the FK-Ising model and the Ising model with parameter $p=1-\exp(-\frac{2}{T})$ can be described as follows: for $G=(V,E)$, consider a probability measure $\pi$ on $\Sigma\otimes\Xi$ defined by
$\pi(\sigma,\omega)=\frac{1}{\mathcal{Z}_{\pi}}\prod_{e=\{x,y\}\in E}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\delta(\sigma_{x},\sigma_{y})\Big{]},$
where $\delta(\cdot,\cdot)$ is a $\{0,1\}$-valued function such that $\delta(a,b)=1$ if and only if $a=b$ and $\mathcal{Z}_{\pi}$ is the normalizing constant. Then we have the following properties:
  1. (i)
    Marginal on $\Sigma$: for a fixed $\sigma\in\Sigma$, summing over all $\omega\in\Xi$ we have
    $\pi(\sigma)=\displaystyle\sum_{\omega\in\Xi}\pi(\sigma,\omega)=\dfrac{1}{\mathcal{Z}_{\pi}}\prod_{e=\{x,y\}\in E}\Big{[}(1-p)+p\delta(\sigma_{x},\sigma_{y})\Big{]}\propto(1-p)^{\sum_{\{x,y\}\in E}\mathbf{1}_{\{\sigma_{x}\neq\sigma_{y}\}}}.$
    This coincides with the Ising measure $\mu_{T,G}^{0,0}$ with 00 boundary condition and $1-p=\exp(-\frac{2}{T})$.
  2. (ii)
    Marginal on $\Xi$: for a fixed $\omega\in\Xi$, summing over all $\sigma\in\Sigma$ gives
    $\pi(\omega)=\sum_{\sigma\in\Sigma}\pi(\sigma,\omega)=\dfrac{1}{\mathcal{Z}_{\pi}}\prod_{\omega(e)=0}(1-p)\sum_{\sigma\in\Sigma}\prod_{\omega(e)=1}p\delta(\sigma_{x},\sigma_{y}).$
    In order for the preceding product to be non-zero, the two spins on the endpoints of every open edge must agree and thus the spins on every open cluster are the same. As a result, there are $2^{\kappa(\omega)}$ different choices for $\sigma$, thus we have
    $\pi(\omega)=\dfrac{1}{\mathcal{Z}_{\pi}}\prod_{e\in E}\Big{[}p^{\omega(e)}(1-p)^{1-\omega(e)}\Big{]}\times 2^{\kappa(\omega)},$
    which corresponds to the measure $\phi^{\mathrm{f},0}_{p,G}$ of the FK-Ising model on $\{0,1\}^{E}$.
  3. (iii)
    Conditioned on a spin configuration: for a fixed $\sigma\in\Sigma$, let $E_{1}=\{\{x,y\}\in E:\sigma_{x}=\sigma_{y}\}$ and $E_{2}=E\setminus E_{1}$. Dividing $\pi(\sigma,\omega)$ by $\pi(\sigma)$ yields that
    $\pi_{|\sigma}(\omega)\propto\prod_{e\in E_{2}}\delta(\omega(e),0)\prod_{e\in E_{1}}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\Big{]}.$
    Thus, under $\pi_{|\sigma}$, the edges in $E_{2}$ must be closed and in addition, edges in $E_{1}$ are open with probability $p$ and closed with probability $1-p$ independently. In other words, the conditional measure for $\omega$ can be viewed as a Bernoulli bond percolation within each spin cluster.
  4. (iv)
    Conditioned on an edge configuration: for a fixed $\omega\in\Xi$ let $E_{3}=\{e\in E:\omega(e)=1\}$ and $E_{4}=E\setminus E_{3}$. Then we have
    $\pi_{|\omega}(\sigma)\propto\prod_{e\in\{x,y\}\in E_{3}}p\delta(\sigma_{x},\sigma_{y}).$
    Thus, the conditional measure $\pi_{|\omega}$ is uniform among spin configurations where each open cluster in $\omega$ receives the same spin.
Furthermore, the above coupling can be extended to the case with external field $\{\epsilon h_{x}:x\in V\}$. To this end, we define
$\pi_{h}(\sigma,\omega)=\frac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e=\{x,y\}\in E}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\delta(\sigma_{x},\sigma_{y})\Big{]}\cdot\exp\left(-\frac{1}{T}\sum_{x\in V}\epsilon h_{x}\sigma_{x}\right),$
where $\mathcal{Z}_{\pi_{h}}$ is the normalizing constant. By a straightforward computation, we see
$\mathcal{Z}_{\pi_{h}}=\sum_{\sigma\in\Sigma}\sum_{\omega\in\Xi}\pi_{h}(\sigma,\omega)=\sum_{\sigma\in\Sigma}\exp\left[\frac{1}{T}\left(\sum_{\{x,y\}\in E}\sigma_{x}\sigma_{y}+\sum_{x\in V}\epsilon h_{x}\sigma_{x}\right)\right]\cdot\exp\left(-\frac{1}{T}|E|\right),$
implying that the normalizing constant $\mathcal{Z}_{\pi_{h}}$ and the partition function $\mathcal{Z}_{\mu_{h}}$ for the Ising measure $\mu_{T,G}^{\xi,\epsilon h}$ are equal up to a factor of $\exp(-\frac{1}{T}|E|)$, which in particular does not depend on the external field $\epsilon h$. One can verify that the aforementioned properties (in the case without external field) can be extended as follows:
  1. (i)
    For a fixed $\sigma\in\Sigma$, the additional term $\exp(\frac{1}{T}\sum_{x\in V}\epsilon h_{x}\sigma_{x})$ is a constant in $\omega$. As a result, the marginal distribution of $\pi_{h}$ on $\Sigma$ is exactly the Ising measure with external field $\epsilon h$. In addition, the conditional distribution for percolation configuration given a fixed $\sigma$ is also equivalent to a Bernoulli bond percolation within each spin cluster.
  2. (ii)
    For a fixed $\omega\in\Xi$, in order for $\pi_{h}(\sigma,\omega)$ to be non-zero, each open cluster must have the same spin. Thus, writing $h_{A}=\sum_{x\in A}h_{x}$ for any subset $A\subset V$, we have
    $\sum_{\sigma\in\Sigma}\pi_{h}(\sigma,\omega)=\dfrac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e\in E}p^{\omega(e)}(1-p)^{1-\omega(e)}\prod_{j=1}^{\kappa(\omega)}2\cosh(h_{\mathcal{C}_{j}}/T).$
    Therefore, conditioned on a fixed $\omega\in\Xi$, each open cluster $\mathcal{C}_{j}$ must have the same spin and this spin is plus with probability $\frac{\exp(h_{\mathcal{C}_{j}}/T)}{2\cosh(h_{\mathcal{C}_{j}}/T)}$.
[…]
2.5 Basic properties of the model(s)
Here we state some of the standard and well-known properties that will be used repeatedly.
  1. (i)
    FKG inequality. This was introduced in [24] and named after the three authors.
    If $A,B$ are both increasing (or decreasing) events, then we have
    $P(A\cap B)\geq P(A)\times P(B).$
  2. (ii)
    Comparison of boundary conditions. If two boundary conditions $\xi_{1}\leq\xi_{2}$, then for any increasing event $A$ we have
    $P^{\xi_{1}}(A)\leq P^{\xi_{2}}(A).$
    Here $\xi_{1}\leq\xi_{2}$ stands for a partial order in the set of boundary conditions: for spin configurations, we say $\xi_{1}\leq\xi_{2}$ if every plus spin in $\xi_{1}$ is also plus in $\xi_{2}$; for edge configurations, we say $\xi_{1}\leq\xi_{2}$ if every open edge in $\xi_{1}$ is also open in $\xi_{2}$.
  3. (iii)
    Domain Markov property. For two domains $\Gamma_{1}\subset\Gamma_{2}$, given the configuration $\xi$ on $\Gamma_{2}/\Gamma_{1}$, the influence on the measure in $\Gamma_{1}$ behaves like a boundary condition:
    $P_{\Gamma_{2}}(\,\cdot\,|\xi)=P_{\Gamma_{1}}^{\xi_{|\partial\Gamma_{1}}}(\cdot),$
    where on the right-hand side $\xi_{|\partial\Gamma_{1}}$ is the restriction on $\partial_{ext}\Gamma_{1}$ if $P$ is an Ising measure and $\xi_{|\partial\Gamma_{1}}$ is a partition on $\partial_{int}\Gamma_{1}$ if $P$ is an FK-Ising measure.
  4. (iv)
    Dual path. We review the dual theory for the FK-Ising model initiated in [30], and our presentation follows that in [4]. Define $(\mathbb{Z}^{2})^{\diamond}=\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})$ to be the dual of $\mathbb{Z}^{2}$, which can be viewed as a translation of $\mathbb{Z}^{2}$ in $\mathbb{R}^{2}$. We see that every vertex in $(\mathbb{Z}^{2})^{\diamond}$ is the center of a unit square in $\mathbb{Z}^{2}$ and vice versa. In other words, every edge $e\in\mathbb{Z}^{2}$ intersects with a unique edge $e^{\diamond}\in(\mathbb{Z}^{2})^{\diamond}$. For a configuration $\omega\in\{0,1\}^{E(\mathbb{Z}^{2})}$, we define its dual configuration $\omega^{\diamond}$ by setting
    $\omega^{\diamond}(e^{\diamond})=1-\omega(e),~{}~{}~{}\mbox{for~{}all~{}}e\in E(\mathbb{Z}^{2}).$
    Importantly, the measure of $\omega^{\diamond}$ is also an FK-Ising measure with a dual boundary condition and dual parameter $p^{\diamond}$ (see [4] for more details), whereby [35]
    $p_{c}^{\diamond}=p_{c}=\dfrac{\sqrt{2}}{1+\sqrt{2}}.$
[…]
In (i), (ii), and (iii), $P$ may stand for the Ising measure with or without the external field and with or without the boundary condition and may also stand for the FK-Ising measure without the external field. However, the FK-Ising measure with the external field only satisfies (i) and (ii). We will write FKG, CBC, and DMP to represent the properties (i), (ii), and (iii) in the following for convenience.
2.6 Boundary influence and correlation length
One of the key quantities of interest for the RFIM is the boundary influence defined as follows:
$m(T,N,\epsilon):=\mathbb{E}(\tilde{m}(T,N,\epsilon h))=\frac{1}{2}\mathbb{E}\left(\langle\sigma_{o}\rangle^{+,\epsilon h}_{T,\Lambda_{N}}-\langle\sigma_{o}\rangle^{-,\epsilon h}_{T,\Lambda_{N}}\right),$
where $\sigma\in\{-1,1\}^{\Lambda_{N}}$ is the spin configuration, $\epsilon h=\{\epsilon h_{v}:v\in\mathbb{Z}^{d}\}$ is the external field with strength $\epsilon>0$. And we will say the long-range order exists for the RFIM if the boundary influence has a positive lower bound as the box size $N$ goes to infinity. The absence of long-range order in two dimensions for any temperature $T$ and any disorder strength $\epsilon>0$ raises an intriguing question: can we still observe a phase transition phenomenon for the 2D RFIM? The answer is positive, the behavior of boundary influence alters when $\epsilon$ decays with $N$ in different rates.
More precisely, Define the correlation length of two-dimensional RFIM as follows:
${\psi_{\star}(T,\epsilon)}=\min\{N:m(T,N,\epsilon)\leq m(T,N,0)/2\}\,.$
By a sequence of works [18, 20, 15], they prove that the correlation length has order $ce^{c\epsilon^{-4/3}}$ at low temperature $T<T_{c}$ and has order $c\epsilon^{-8/7}$ at critical temperature $T=T_{c}$. Furthermore, [15] proves when $T=T_{c}(2)$:
  • For $\epsilon\ll N^{-\frac{7}{8}},$ we have $m(T,N,\epsilon)=(1+o(1))m(T,N,0).$
  • For $\epsilon\gg N^{-\frac{7}{8}}$, the boundary influence $m(T,N,\epsilon)$ decays exponentially in $N$, moreover, the decay rate is of order $\epsilon^{\frac{8}{7}}.$
In summary, at the critical temperature, the boundary influence can “perceive” the random external field if and only if $\epsilon\gg N^{-7/8}$. In contrast to this threshold, our result Theorem 1.1 implies that for any statistic of the Ising measure at any temperature, as long as $\epsilon\ll N^{-1}$, it will behave similarly in both the cases with and without random field. Furthermore, Theorem 1.3 implies that for any statistic of the FK-Ising measure at the critical-temperature, as long as $\epsilon\ll N^{-\frac{15}{16}}$, it will behave similarly in both the cases with and without random field.
3 Proof of Theorem 1.1
We begin by proving Theorem 1.1. While its proof is significantly less complex than those of Theorem 1.3, it introduces key concepts that we will repeatedly utilize throughout our discussion.
We will refer to inequalities such as (1.1) as “absolute continuity” and (1.2) as “singularity” for the sake of simplicity, although this terminology is not mathematically rigorous since all the measures discussed here are discrete. Notice that the total variation distance is always between 0 and 1, thus we can always assume $\epsilon N$ is small enough while proving (1.1) and large enough while proving (1.2).
3.1 Obtaining absolute continuity from decoupling
It is easy to calculate that
$\frac{\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}(\sigma)}{\mu^{\xi,0}_{T,\Lambda_{N}}(\sigma)}=\frac{\mathcal{Z}^{\xi,0}_{\mu,T,\Lambda_{N}}}{\mathcal{Z}^{\xi,\epsilon h}_{\mu,T,\Lambda_{N}}}\exp\left(\epsilon\sum_{x\in\Lambda_{N}}\sigma_{x}h_{x}\right).$ (3.1)
For any $\sigma$ fixed, we have that $\epsilon\sum_{x\in\Lambda_{N}}\sigma_{x}h_{x}$ is a Gaussian variable with variance $\epsilon^{2}N^{2}$, therefore it concentrates well when $\epsilon N$ is sufficiently small. But we have to deal with all possible $\sigma$ for a fixed external field $\epsilon h.$ To address this, we have the following decoupling lemma:
Lemma 3.1.
Let $(\Omega,\mathcal{F})$ be a measurable space and $\nu$ is a probability measure on it. Let $\eta$ be a random field with law $\mathbb{Q}$ and $\nu^{\eta}$ is a random probability measure on $(\Omega,\mathcal{F})$. For any $\omega\in\Xi,$ define $G(\eta,\omega)=\frac{\nu^{\eta}(\omega)}{\nu(\omega)},$ then we have for any constants $10<a<b<$ 1:
$\mathbb{Q}(\|\nu^{\eta}-\nu\|_{\mathrm{TV}}\geq b)\leq\frac{\mathbb{Q}\otimes\nu\left(G(\eta,\omega)<1-a\right)}{b-a}.$
Proof.
For any fixed external field $\eta$, we can first rewrite the total variation distance as follows:
$\|\nu^{\eta}-\nu\|_{\mathrm{TV}}$ $=\sum_{\omega\in\Xi}\left(\nu(\omega)-\nu^{\eta}(\omega)\right)_{+}=\sum_{\omega\in\Xi}\nu(\omega)\left(1-G(\eta,\omega)\right)_{+}$
$=\sum_{\omega:G(\eta,\omega)\geq 1-a}\nu(\omega)\left(1-G(\eta,\omega)\right)_{+}+\sum_{\omega:G(\eta,\omega)<1-a}\nu(\omega)\left(1-G(\eta,\omega)\right)_{+}$
$\leq a+\nu\left(G(\eta,\omega)<1-a\right),$
where $(x)_{+}=\max\{x,0\}$. Thus, we have
$\mathbb{Q}(\|\nu^{\eta}-\nu\|_{\mathrm{TV}}\geq b)\leq\mathbb{Q}(\nu(G(\eta,\omega)<1-a)\geq b-a)\leq\frac{\mathbb{Q}\otimes\nu\left(G(\eta,\omega)<1-a\right)}{b-a}.\qed$
In a word, by losing acceptable accuracy, we can focus on the product measure $\mathbb{Q}\otimes\nu$ instead of the random measure $\nu^{\eta}.$ This is sufficient to prove the absolute continuity.
Proof of (1.1).
Taking $\nu=\mu^{\xi,0}_{T,\Lambda_{N}}$ and $\nu^{\eta}=\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}$ in Lemma 3.1, all we need is to upper-bound
$G(h,\sigma)=\frac{\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}(\sigma)}{\mu^{\xi,0}_{T,\Lambda_{N}}(\sigma)}=\frac{\mathcal{Z}^{\xi,0}_{\mu,T,\Lambda_{N}}}{\mathcal{Z}^{\xi,\epsilon h}_{\mu,T,\Lambda_{N}}}\exp\left(\epsilon\sum_{x\in\Lambda_{N}}\sigma_{x}h_{x}\right).$
Write
$Z(h)=\frac{\mathcal{Z}^{\xi,0}_{\mu,T,\Lambda_{N}}}{\mathcal{Z}^{\xi,\epsilon h}_{\mu,T,\Lambda_{N}}}\text{ and }E(h,\sigma)=\exp\left(\epsilon\sum_{x\in\Lambda_{N}}\sigma_{x}h_{x}\right)$
for short. Then by chaos expansion technique ([12, Theorem 3.14]) for Ising (see Section 4.1 for a detailed proof), we have
$\mathbb{P}\left(|Z(h)-1|\leq\sqrt{\epsilon N}\right)\geq 1-c_{1}\exp\left(-c_{1}\sqrt{\epsilon^{-1}N^{-1}}\right).$
Meanwhile, for any fixed $\sigma,$ we have $\epsilon\underset{{}_{x\in\Lambda_{N}}}{\sum}\sigma_{x}h_{x}$ is a Gaussian variable with variance $\epsilon^{2}N^{2}.$ Therefore,
$\mathbb{P}\left(E(h,\sigma)\leq 1-c_{2}\sqrt{\epsilon N}\right)\leq c_{2}\exp\left(-c_{2}^{-1}\epsilon^{-1}N^{-1}\right).$
Thus, we have
$\mathbb{P}\otimes\mu^{{\xi,0}}_{{T,\Lambda_{N}}}\left(G(h,\sigma)\leq 1-c_{3}\sqrt{\epsilon N}\right)\leq c_{3}\exp\left(-c_{3}\sqrt{\epsilon^{-1}N^{-1}}\right),$
by choosing $a=c_{3}\sqrt{\epsilon N}<b=c_{3}(\epsilon N)^{1/4}$ in Lemma 3.1 and recalling $\epsilon N$ is small enough, the proof is completed. ∎
3.2 Obtaining singularity from independency
The idea behind proving singularity for the RFIM primarily involves leveraging the independence obtained from the DMP. First, we have the following natural result for the total variation distances between product measures.
Definition 3.2.
Let $\{\mu_{i}:i=1,2,\cdots,n\}$ be a sequence of distributions on measurable spaces $\{(\mathcal{E}_{i},\mathcal{F}_{i}):i=1,2,\cdots,n\}.$ Then for any event $\mathcal{A}_{i}\in\mathcal{F}_{i}$, we define $\hat{\mathcal{A}}_{i}=\mathcal{E}_{1}\otimes\cdots\otimes\mathcal{E}_{i-1}\otimes\mathcal{A}_{i}\otimes\mathcal{E}_{i+1}\otimes\cdots\otimes\mathcal{E}_{n}$ to be a cylinder set.
Lemma 3.3.
Let $\mu_{1},\mu_{2},\cdots,\mu_{n}$ and $\nu_{1},\nu_{2},\cdots,\nu_{n}$ be two sequences of distributions with $\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a>0$ and $\mu=\otimes_{i=1}^{n}\mu_{i},~{}\nu=\otimes_{i=1}^{n}\nu_{i}$ be their product measure. Then there exists an absolute constant $c_{1}>0$ such that
$\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp\left(-{c_{1}}^{-1}a^{2}n\right).$
Proof.
Since $\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a$, there exists a event $\mathcal{A}_{i}\in\mathcal{F}_{i}$ such that $\mu_{i}(\mathcal{A}_{i})-\nu_{i}(\mathcal{A}_{i})\geq a$. Since $\mathcal{A}_{i}(i=1,2,\cdots,n)$ are independent under $\mu$, by Lemma B.6, we obtain that
$\mu\left(\Big{|}\sum_{i=1}^{n}\mathbf{1}_{\hat{\mathcal{A}}_{i}}-\mu(\hat{\mathcal{A}}_{i})\Big{|}{\geq}\frac{na}{2}\right)$ $\leq 2\exp\left(-C_{1}a^{2}n\right),$ (3.2)
$\nu\left(\Big{|}\sum_{i=1}^{n}\mathbf{1}_{\hat{\mathcal{A}}_{i}}-\nu(\hat{\mathcal{A}}_{i})\Big{|}{\geq}\frac{na}{2}\right)$ $\leq 2\exp\left(-C_{1}a^{2}n\right).$ (3.3)
Since $\mu(\hat{\mathcal{A}}_{i})-\nu(\hat{\mathcal{A}}_{i})=\mu_{i}(\mathcal{A}_{i})-\nu_{i}(\mathcal{A}_{i})\geq a$, we get that
$\left\{\Big{|}\sum_{i=1}^{n}\mathbf{1}_{\hat{\mathcal{A}}_{i}}-\mu(\hat{\mathcal{A}}_{i})\Big{|}{<}\frac{na}{2}\right\}\cap\left\{\Big{|}\sum_{i=1}^{n}\mathbf{1}_{\hat{\mathcal{A}}_{i}}-\nu(\hat{\mathcal{A}}_{i})\Big{|}{<}\frac{na}{2}\right\}=\emptyset.$
Combined with (3.2) and (3.3), it yields that $\|\mu-\nu\|_{\mathrm{TV}}\geq 1-2\exp(-C_{1}a^{2}n)$. ∎
Corollary 3.4.
Let $\mu_{1},\mu_{2},\cdots,\mu_{n}$ and $\nu_{1},\nu_{2},\cdots,\nu_{n}$ be two sequences of independent random distributions with joint law $\mathbb{P}$. Assume ${\mathbb{E}}\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a>0$ for all $1\leq i\leq n$ where $\mathbb{E}$ is the expectation w.r.t. $\mathbb{P}$, and $\mu=\otimes_{i=1}^{n}\mu_{i},~{}\nu=\otimes_{i=1}^{n}\nu_{i}$ are product measures. Then there exists an absolute constant $c_{1}>0$ such that
$\mathbb{P}\Big{(}\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp(-c_{1}^{-1}{a^{2}}n)\Big{)}\geq 1-c_{1}\exp\left({-}c_{1}^{-1}a^{2}n\right).$
Proof.
Let $I\subset\{1,2,\cdots,n\}$ denote the random set of indices such that $\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq\frac{a}{2}$. Since ${\mathbb{E}}\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a$, we get that $\mathbb{P}(\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq\frac{a}{2})\geq\frac{a}{2}$. Combined with the fact that $\{\mu_{i}\}_{1\leq i\leq n}$ and $\{\nu_{i}\}_{1\leq i\leq n}$ are two independent sequences of distributions, it yields by Lemma B.6 that
$\mathbb{P}\left(|I|\geq\frac{an}{4}\right)\geq 1-C_{1}\exp(-C_{1}^{-1}a^{2}n).$ (3.4)
The rest of the proof is the same as that of Lemma 3.3. ∎
Now we are ready to prove the singularity result.
Proof of (1.2).
Write
$U_{N}=(2\mathbb{Z})^{2}\cap\Lambda_{N},\quad V_{N}=\Lambda_{N}\setminus U_{N};\text{ and }\Sigma_{U}=\{-1,1\}^{U_{N}},\Sigma_{V}=\{-1,1\}^{V_{N}}.$
Thus, for any $x\in U_{N},$ all the neighboring points of $x$ are contained in $V_{N}.$ For any $\sigma_{U}\in\Sigma_{U}$ and $\sigma_{V}\in\Sigma_{V}$, we will write $\sigma=\sigma_{U}\oplus\sigma_{V}$ for $\sigma\in\Sigma=\{-1,1\}^{\Lambda_{N}}$ if the restriction of $\sigma$ on $U_{N}$ and $V_{N}$ coincides with $\sigma_{U}$ and $\sigma_{V}$ respectively. Then we can decompose the total variation distance as follows:
$1-\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}=\sum_{\sigma\in\Sigma}\min\left\{\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}(\sigma),\mu^{\xi,0}_{T,\Lambda_{N}}(\sigma)\right\}$
[…]
Note that we can first couple the configuration on $V_{N}$ and throw the cases that the coupling failed, thus we get
$\min\left\{\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}(\sigma_{U}\oplus\sigma_{V}),\mu^{\xi,0}_{T,\Lambda_{N}}(\sigma_{U}\oplus\sigma_{V})\right\}$
[…]
Combined with (3.5), it yields that
$1-\|\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}-\mu^{\xi,0}_{T,\Lambda_{N}}\|_{\mathrm{TV}}$
[…]
By DMP, the distribution of points in $U_{N}$ are independent of each other. Let $V_{x}$ denote the set of neighbours of $x$, and write $\sigma_{V_{x}}=(\sigma_{V})_{\mid V_{x}}$ then we have
$\mu_{T,U_{N}}^{\sigma_{V},\epsilon h}=\underset{x\in U_{N}}{\otimes}\mu_{T,\{x\}}^{\sigma_{V_{x}},\epsilon h_{x}}~{}~{}\mbox{and}~{}~{}\mu_{T,U_{N}}^{\sigma_{V},0}=\underset{x\in U_{N}}{\otimes}\mu_{T,\{x\}}^{\sigma_{V_{x}},0}.$
Note that
$\Big{|}\mu_{T,\{x\}}^{\sigma_{V_{x}},\epsilon h_{x}}(\sigma_{x}=1)-\mu_{T,\{x\}}^{\sigma_{V_{x}},0}(\sigma_{x}=1)\Big{|}=\Big{|}g\left(\frac{N_{x}}{T}\right)-g\left(\frac{N_{x}+\epsilon h_{x}}{T}\right)\Big{|}=\Big{|}\frac{\epsilon h_{x}}{T}g^{\prime}\left(\frac{N_{x}}{T}+\theta\right)\Big{|},$
where we write $N_{x}=\sum_{y\sim x}\sigma_{V_{x}}(y)$, $g(t)=\frac{e^{t}}{e^{t}+e^{-t}},$ and $\theta$ is a constant between 0 and $\frac{\epsilon h_{x}}{T}$. Since $|g^{\prime}(t)|$ decays in $|t|$ and $|N_{x}|\leq 4,$ we have
$\|\mu_{T,\{x\}}^{\sigma_{V_{x}},\epsilon h_{x}}-\mu_{T,\{x\}}^{\sigma_{V_{x}},0}\|_{\mathrm{TV}}\geq\frac{\epsilon|h_{x}|}{4T}e^{-4-\epsilon|h_{x}|}.$
Combined with Corollary 3.4 and the fact that $|h_{x}|\in[1,2]$ with positive probability, it yields the desired result. ∎
4 Absolute continuity in weak-disorder regime
From now on, we focus on the FK-Ising model and the random field FK-Ising model. In this section, we will prove the absolute continuity results for all temperature regimes. The proof is similar to that in Section 3.1, but some new ideas are required. Since many calculations in high-temperature, low-temperature, and critical-temperature regimes are similar, we introduce the following notation for simplicity:
$\beta(T)=\left\{\begin{aligned} &1~{}~{}~{}&0<T<T_{c}\,,\\ &\frac{7}{8}~{}~{}~{}&T=T_{c}\,,\\ &\frac{1}{2}~{}~{}~{}&T>T_{c}\,.\end{aligned}\right.$
Our main goal is to show that
$\frac{\phi^{\gamma,\epsilon h}_{p,\Lambda_{N}}(\omega)}{\phi^{\gamma,0}_{p,\Lambda_{N}}(\omega)}=Z(h)\exp\left(F(h,\omega)\right)$ (4.1)
is close to 1 for typical disorder when $\epsilon\ll N^{-\beta(T)}$ where
$Z(h):=\frac{\mathcal{Z}^{\gamma,0}_{\phi,p,\Lambda_{N}}}{\mathcal{Z}^{\gamma,\epsilon h}_{\phi,p,\Lambda_{N}}},\quad F(h,\omega)=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}f{\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)}.$ (4.2)
In Section 4.1, we show the concentration of $Z(h)$ using a similar technique as in [12, Theorem 3.14] and in Section 4.2, we apply Lemma 3.1 to show the absolute continuity.
4.1 Concentration of $Z(h)$
We start by controlling the ratio of partition functions with and without external field. We expand the ratio between the partition functions with and without disorder as follows:
$\frac{1}{Z(h)}=\frac{\mathcal{Z}^{\gamma,\epsilon h}_{\phi,p,\Lambda_{N}}}{\mathcal{Z}^{\gamma,0}_{\phi,p,\Lambda_{N}}}=\frac{\mathcal{Z}^{\gamma,\epsilon h}_{\mu,T,\Lambda_{N}}}{\mathcal{Z}^{\gamma,0}_{\mu,T,\Lambda_{N}}}$ $=\Bigg{\langle}~{}\exp\left(\sum_{v\in\Lambda_{N}}\frac{\epsilon h_{v}\sigma_{v}}{T}\right)~{}\Bigg{\rangle}_{T,\Lambda_{N}}^{\gamma,0}$
$=\Bigg{\langle}~{}\prod_{v\in\Lambda_{N}}\left(\cosh\left(\frac{\epsilon h_{v}}{T}\right)+\sigma_{v}\sinh\left(\frac{\epsilon h_{v}}{T}\right)\right)~{}\Bigg{\rangle}_{T,\Lambda_{N}}^{\gamma,0}$
$=\prod_{v\in\Lambda_{N}}\cosh\left(\frac{\epsilon h_{v}}{T}\right)\Bigg{\langle}~{}\prod_{v\in\Lambda_{N}}\left(1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\right)~{}\Bigg{\rangle}_{T,\Lambda_{N}}^{\gamma,0}.$ (4.3)
To control the right-hand side of (4.3), we recall the following techniques from [6] and [15].
Definition 4.1.
Let $\Omega$ be a region such that $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and $\gamma\in\{-1,1\}^{\partial_{int}\Omega}$ be a boundary condition on $\partial_{int}\Omega$. Let $\mathcal{A}_{0}\subset{\mathbb{R}}^{\Omega}\otimes\Xi$ denote the collection of pairs $(h,\omega)$ such that the property (P0) holds,
$(\text{\emph{P}}0):\quad\Bigg{|}\Bigg{\langle}~{}\prod_{v\in\Omega}\Bigg{[}1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\Bigg{]}~{}\Bigg{\rangle}^{\gamma,0}_{T,\Omega}-1\Bigg{|}\leq\sqrt{\epsilon N^{\beta(T)}}.$ (4.4)
Note that actually $\mathcal{A}_{0}$ depends only on $h$, so we could define $\mathcal{A}_{0}\subset{\mathbb{R}}^{\Omega}$ to be the collection of the external field $h$ such that the property (P0) holds. However, for further convenience, we keep $\omega$ in the definition of $\mathcal{A}_{0}$.
Remark 4.2.
Of course, the set $\mathcal{A}_{0}$ relies on $\gamma,\Omega,T$ and $\epsilon,$ but we omit them from the notation for concision. The same conventions will be used several times without further notice.
Lemma 4.3.
Take $\Omega=\Lambda_{N},$ then there exist constants $c_{1},c_{2}>0$ depending only on $T$ such that $\mathbb{P}\otimes\phi_{p,\Omega}^{\gamma,0}(\mathcal{A}^{0})\geq 1-c_{1}\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}N^{-\beta(T)}}\right)$ for any disorder strength $\epsilon\leq c_{2}N^{-\beta(T)}$ and any boundary condition $\gamma$ on $\partial_{int}\Lambda_{N}$.
Proof.
The case $T=T_{c}$ is covered by [15, Lemma 3.4]. The cases that $T<T_{c}$ and $T>T_{c}$ follow a similar proof strategy. Here, we only highlight the key adaptations required. Note that
$\Bigg{\langle}~{}\prod_{v\in\Lambda_{N}}\Bigg{[}1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\Bigg{]}~{}\Bigg{\rangle}^{\gamma,0}_{T,\Lambda_{N}}-1=\sum_{k=1}^{\infty}\sum_{J\subset\Lambda_{N},|J|=k}\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\prod_{v\in J}\tanh\left(\frac{\epsilon h_{v}}{T}\right).$
To control the right-hand side, using [15, Lemma B.1], the core is to bound
$A_{k}=\epsilon\Big{[}\sum_{J\subset\Lambda_{N},|J|=k}\left(\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\right)^{2}\Big{]}^{\frac{1}{2k}}.$
For $T<T_{c}$, using $|\sigma_{x}|=1$, we can obtain
$A_{k}\leq\epsilon\binom{N^{2}}{k}^{\frac{1}{2k}}\leq c_{2}N^{-1}\left(\frac{N^{2k}}{k!}\right)^{\frac{1}{2k}}=\frac{c_{2}}{(k!)^{\frac{1}{2k}}}.$
For $T>T_{c}$, by the Edward-Sokal coupling between the Ising model and the FK-Ising model, we have $\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}=\phi^{\gamma,0}_{p,\Lambda_{N}}(\mathtt{Even}_{J}),$ where $\mathtt{Even}_{J}$ stands for the event that each cluster in $\omega^{\gamma}$ contains an even number of points in $J$.
Furthermore, the probability of a cluster being large decays exponentially when $p<p_{c},$ i.e., when $T>T_{c}$, we have for any boundary condition $\gamma$ (see for example [22]):
$\phi_{p,\Omega}^{\gamma,0}\left(x\longleftrightarrow\partial_{int}\Lambda_{m}(x)\right)\leq e^{-cm}.$ (4.5)
For any $J\subset\Lambda_{N}$, let $d_{x}=\frac{1}{2}\cdot dist(x,\partial_{int}\Lambda_{N}\cup J\setminus\{x\})$ and let $\mathcal{R}_{x}=\Lambda_{d_{x}}(x)$. Then it is easy to get that
$\mathtt{Even}_{J}\subset\bigcap_{x\in J}\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}.$ (4.6)
Combining (4.5) and (4.6), we get that
$\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\leq\prod_{x\in J}\phi_{p,\mathcal{R}_{x}}^{\mathrm{w},0}\left(\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}\right)\leq\exp{\Big{(}-\sum_{x\in J}c\cdot d_{x}\Big{)}}.$ (4.7)
Then we get from the calculation in [12, Lemmas 8.1 and 8.3] that
$\sum_{J\subset\Lambda_{N},|J|=k}\left(\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\right)^{2}\leq\frac{\prod_{j=1}^{k}(C_{1}N+C_{1}k)}{k!}.$ (4.8)
Thus, we get that
$A_{k}$ $\leq\epsilon\left(\frac{\prod_{j=1}^{k}(C_{1}N+C_{1}k)}{k!}\right)^{\frac{1}{2k}}\leq\epsilon\left(\frac{(C_{1}N+C_{1}k)^{k}}{k!}\right)^{\frac{1}{2k}}$
$\leq\epsilon\cdot\frac{\sqrt{C_{1}N+C_{1}k}}{\sqrt{k/e}}\leq\frac{\sqrt{C_{1}N}\epsilon}{\sqrt{k/e}}+\sqrt{eC_{1}}\epsilon.\qed$
With the upper bound for $A_{k}$, the rest of the proof is the same to that in [15, Lemma 3.4].
In order to meet further needs, we prove a stronger version of Lemma 4.3 even though it will not be used in this section.
Lemma 4.4.
Let $\Omega$ be a region such that $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and $\gamma\in\{-1,1\}^{\partial_{int}\Omega}$ be a boundary condition on $\partial_{int}\Omega$.
  1. 1.
    If $T\geq T_{c}$ and $\gamma=\mathrm{f}$ is the free boundary condition, then there exist constants $c_{1},c_{2}>0$ depending only on $T$ such that $\mathbb{P}(\mathcal{H}^{0})\geq 1-c_{1}\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}N^{-\beta(T)}}\right)$ for all disorder strength $\epsilon\leq c_{2}N^{-\beta(T)}$.
  2. 2.
    If $T<T_{c}$, then there exist constants $c_{3},c_{4}>0$ depending only on $T$ such that $\mathbb{P}(\mathcal{H}^{0})\geq 1-c_{3}\exp\left(-c_{3}^{-1}\sqrt{\epsilon^{-1}N^{-\beta(T)}}\right)$ for all disorder strength $\epsilon\leq c_{4}N^{-\beta(T)}$ and any boundary condition $\gamma$.
Proof.
The proof is similar to the proof of Lemma 4.3, so we follow the notations there. The case that $T<T_{c}$ follows directly from the proof of Lemma 4.3.
If $T\geq T_{c}$ and $\gamma=\mathrm{f}$ is the free boundary, then we get from CBC that $\langle\sigma^{J}\rangle^{\mathrm{f},0}_{T,\Omega}\leq\langle\sigma^{J}\rangle^{\mathrm{w},0}_{T,\Lambda_{2N}}.$ The desired bound on $A_{k}$ then follows from the proof of Lemma 4.3. ∎
Further, in order to control the product of hyperbolic cosine in the right-hand side of (4.3), we give the following definition.
Definition 4.5.
Let $\mathcal{A}_{1}\subset{\mathbb{R}}^{\Omega}\otimes\Xi$ denote the collection of pairs $(h,\omega)$ such that the property (P1) holds
$(\text{\emph{P}}1):\quad\Big{|}\sum_{v\in\Lambda_{N}}f\left(\frac{\epsilon h_{v}}{T}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\Big{|}\leq\sqrt{\epsilon^{2}N}.$ (4.9)
Lemma 4.6.
There exist constants $c_{1}>0$ and $c_{2}>0$ depending only on $T$ such that for any disorder strength $\epsilon\leq c_{1}N^{-\frac{1}{2}}$, we have $\mathbb{P}\otimes\phi_{p,\Omega}^{\gamma,0}(\mathcal{A}_{1})\geq 1-c_{2}\exp\left(-c_{2}^{-1}\epsilon^{-1}N^{-\frac{1}{2}}\right)$.
Proof.
Recall that $f(x)=\ln\cosh(x)$. Given the inequality $\frac{x^{2}-x^{4}}{2}\leq f(x)\leq\frac{x^{2}}{2}$, we calculate that
$\mathbb{P}\left(\Big{|}\sum_{v\in\Lambda_{N}}f\left(\frac{\epsilon h_{v}}{T}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\Big{|}>\sqrt{\epsilon^{2}N}\right)$
[…]
where the last inequality comes from Lemma B.5 and letting $c_{1}>0$ small enough. ∎
Combining (4.3), (4.4) and (4.9), we obtain that for $(h,\omega)\in\mathcal{A}_{0}\cap\mathcal{A}_{1},$
$\left(1-\sqrt{\epsilon N^{\beta(T)}}\right)\left(1-\sqrt{\epsilon^{2}N}\right)$ $\leq Z(h)\cdot\exp\left(\frac{2\epsilon^{2}N^{2}}{T^{2}}\right)$ $\leq\left(1+\sqrt{\epsilon N^{\beta(T)}}\right)\left(1+\sqrt{\epsilon^{2}N}\right).$ (4.11)
4.2 Concentration of $F(h,\omega)$
Recalling the definition of $F(h,\omega)$ in (4.2) and noting that $\frac{x^{2}-x^{4}}{2}\leq f(x)\leq\frac{x^{2}}{2}$, we have
$\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{4}h_{\mathcal{C}}^{4}}{2T^{4}}\leq F(h,\omega)\leq\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}.$ (4.12)
Combined with Lemma 3.1, we can consider the “typical” event such that $F(h,\omega)$ concentrates well under the measure $\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}$.
Definition 4.7.
We denote by $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ the collections of pairs $(h,\omega)$ such that the following properties (P2) and (P3) hold respectively.
$(\text{\emph{P}}2):\quad F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\leq\epsilon N^{\alpha(T)},$ (4.13)
$(\text{\emph{P}}3):\quad F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\geq-\epsilon N^{\alpha(T)}.$ (4.14)
Lemma 4.8.
There exist constants $c_{1}=c_{1}(T)>0$ and $c_{2}=c_{2}(T)>0$ such that for any disorder strength $\epsilon<c_{1}N^{-\alpha(T)}$ we have
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}_{2}\cap\mathcal{A}_{3})\geq 1-c_{2}\exp\big{(}-c_{2}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\big{)}.$
Proof.
For any $\omega$ fixed, we write $X_{\mathcal{C}}=\frac{h_{\mathcal{C}}}{\sqrt{|\mathcal{C}|}}$, then we get that $X_{\mathcal{C}}~{}(\mathcal{C}\in\mathfrak{C})$’s are i.i.d. Gaussian variables with mean 0 and variance 1 and we denote the law as $\mathbb{P}_{X}$. Now we can rewrite the second moment term in (4.12) as
$\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)+\frac{2\epsilon^{2}N^{2}}{T^{2}}.$
The first term $\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)$ is a centralized summation of $\chi$-square distributions which can be estimated using Lemma B.5. Let $\mathscr{S}=\{\omega:\sum_{\mathcal{C}\in\mathfrak{C}{(\omega)}}|\mathcal{C}|^{2}\leq\epsilon^{-1}N^{3\alpha(T)}\}.$ Then for any $\omega\in\mathscr{S}$, we have by (4.12)
$\mathbb{P}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}>\epsilon N^{\alpha(T)}\right)$
[…]
The last inequality follows from Lemma B.5. Similarly, we have by (4.12)
$\mathbb{P}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}<-\epsilon N^{\alpha(T)}\right)$
[…]
Again by Lemma B.5, we have
${\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)<-\frac{1}{2}\epsilon N^{\alpha(T)}\right)$ $\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right),$ (4.17)
and $\mbox{and }~{}~{}~{}{\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}}{2T^{4}}X_{\mathcal{C}}^{4}>\frac{1}{2}\epsilon N^{\alpha(T)}\right)$ $\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right).$ (4.18)
Combining (4.17) and (4.18) with (4.16), we get that
$\mathbb{P}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}<-\epsilon N^{\alpha(T)}\right)\leq 2C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right).$ (4.19)
In addition, we have by Corollary C.5 that
$\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathscr{S})\geq 1-C_{2}\exp\left(-C_{2}^{-1}\epsilon^{-1}N^{-\alpha(T)}\right).$ (4.20)
Combining (4.15), (4.19) with (4.20) shows that
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}_{i}^{c})\leq C_{3}\exp\left(-C_{3}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right)$ (4.21)
for $i=2,3,$ some $C_{3}>0$ and sufficiently large $N$. Thus we complete the proof. ∎
Proof of Theorem 1.3: part one.
Let $\mathcal{A}=\mathcal{A}_{0}\cap\mathcal{A}_{1}\cap\mathcal{A}_{2}\cap\mathcal{A}_{3}$, then combining Lemmas 4.3, 4.6 and 4.8 shows
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A})\geq 1-C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right)$
for some constant $C_{1}>0$. For $(h,\omega)\in\mathcal{A}$, by (4.13) and (4.14), we compute that
$\exp(-\epsilon N^{\alpha(T)})\leq\frac{\prod_{\mathcal{C}\in\mathfrak{C}}\cosh(\frac{\epsilon h_{\mathcal{C}}}{T})}{\exp(\frac{2\epsilon^{2}N^{2}}{T^{2}})}\leq\exp\left(\epsilon N^{\alpha(T)}\right).$ (4.22)
Plugging (4.11) and (4.22) into Lemma 3.1 gives the desired result. ∎
5 Near-critical behaviour
The crucial input for Section 6 is to lower-bound the influence of external field on an $M$ box (with $M\sim\epsilon^{-\frac{1}{\alpha(T)}}$) and we will do it in this section for all temperature regimes. To be precise, we will show that as long as $\epsilon>cN^{-\alpha(T)}$ for some constant $c>0$, ${\mathbb{E}}\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{\gamma,\epsilon h}\|_{\mathrm{TV}}$ is bounded away from 00 for $T\geq T_{c}$ in Sections 5.1 and 5.2. A similar lower bound in the low-temperature regime will be obtained in Section 5.3 with different emphasis according to the ingredients that will be used in Section 6.2. Finally, as a byproduct, we will show that $\alpha(T)$ is the critical threshold of the choice of exponent by proving the upper bound for total variation distance on an $M$ box in all temperature regimes in Section 5.4.
We point out that the primary computational effort of this work is concentrated in this section. Readers less interested in technical derivations may proceed directly to Section 6, as the continuity of the exposition remains intact when assuming the core results from Section 5.
5.1 The critical-temperature case
We first consider the critical-temperature case. We prove the following result on general domains that have been used in Section 6.1.
Proposition 5.1.
Fix $T=T_{c}(2)$ and thus $p=p_{c}$. For any constant $\theta>0$, there exists a constant $c=c(\theta)>0$ such that ${\mathbb{E}}\|\phi_{p_{c},\Omega}^{\mathrm{f},0}-\phi_{p_{c},\Omega}^{\mathrm{f},\epsilon h}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-\frac{15}{16}}\leq\epsilon\leq 2\theta N^{-\frac{15}{16}}$, and any domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$.
To obtain the lower bound, recall that in Section 4.1 we have shown that the ratio between the partition functions of the FK-Ising model on the domain $\Omega$ with and without disorder concentrates well. Therefore, it suffices to show that the product of $\cosh(\frac{\epsilon h_{\mathcal{C}}}{T_{c}})$ does not always concentrate at the product of $\exp(\epsilon^{2}|\mathcal{C}|/2T_{c}^{2})$. Moreover, let $\mathcal{C}_{\diamond}$ denote the maximal cluster and note that it has typical volume $cN^{\frac{15}{8}}$, thus we obtain from $\epsilon\geq\theta N^{-\frac{15}{16}}$ that $\epsilon h_{\mathcal{C}_{\diamond}}$ is typically $O(1)$. Hence $\cosh{(\frac{\epsilon h_{\mathcal{C}_{\diamond}}}{T_{c}})}$ will have a positive probability to be much smaller than $\exp(\epsilon^{2}|\mathcal{C}_{\diamond}|/2T_{c}^{2})$, which leads to the anti-concentration result. To carry this out, we introduce the following definition with idea similar to Definition 4.7.
Definition 5.2.
For any constant $c>0$, we use the notation $\mathcal{H}_{\mathrm{cri}}(c)$ to denote the set of external field such that
$\phi_{p_{c},\Omega}^{\mathrm{f},0}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T_{c}}}\right)-\frac{\epsilon^{2}|\Omega|}{2T_{c}^{2}}\leq-1\right)\geq c\,.$
Lemma 5.3.
There exist constants $c_{1},c_{2}>0$ depending only on $\theta$ such that for any disorder strength $\theta N^{-\frac{15}{16}}\leq\epsilon\leq 2\theta N^{-\frac{15}{16}}$, domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and boundary condition $\gamma$ we have $\mathbb{P}(\mathcal{H}_{\mathrm{cri}}(c_{1}))\geq c_{2}$.
Proof.
We still consider the product measure $\mathbb{P}\otimes\phi_{p_{c},\Omega}^{f,0}.$ Let $\mathcal{B}\subset\mathbb{R}^{\Omega}\otimes\{0,1\}^{\mathrm{E}(\Omega)}$ denote the event that
$\sum\limits_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T_{c}}}\right)-\frac{\epsilon^{2}|\Omega|}{2T_{c}^{2}}\leq-1.$ (5.1)
Further, we define $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ to be the following events respectively:
$\sum\limits_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{\diamond}\}}\left(f\left({\frac{\epsilon h_{\mathcal{C}}}{T_{c}}}\right)-\frac{\epsilon^{2}|\mathcal{C}|}{2T_{c}^{2}}\right)\leq 1\mbox{ and }f\left({\frac{\epsilon h_{\mathcal{C}_{\diamond}}}{T_{c}}}\right)-\frac{\epsilon^{2}|\mathcal{C}_{\diamond}|}{2T_{c}^{2}}\leq-2.$
Thus, $\mathcal{B}_{1}\cap\mathcal{B}_{2}\subset\mathcal{B}$. Now we are going to lower-bound the probability of $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. Applying the exponential Markov inequality, we obtain that for any configuration $\omega$
$\mathbb{P}\otimes\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathcal{B}_{1}\mid\omega)\geq 1-e^{-1}\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{\diamond}\}}{\mathbb{E}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T_{c}}\right)\exp\left(\frac{-\epsilon^{2}|\mathcal{C}|}{2T_{c}^{2}}\right)=1-e^{-1}.$ (5.2)
To lower-bound $\mathcal{B}_{2}$, for any configuration $\omega\in\mathscr{S}_{\diamond}=\left\{|\mathcal{C}_{\diamond}|\geq\frac{100T_{c}^{2}}{\epsilon^{2}}\right\}$, we apply the inequality $\cosh(x)\leq\exp(|x|)$ to get that
$\mathbb{P}\otimes\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathcal{B}_{2}\mid\omega)\geq\mathbb{P}_{X}\left(\sqrt{\frac{\epsilon^{2}|\mathcal{C}_{\diamond}|}{T_{c}^{2}}}|X|-\frac{\epsilon^{2}|\mathcal{C}_{\diamond}|}{2T_{c}^{2}}\leq-2\right)\geq 0.9$ (5.3)
where $X$ denotes a normal random variable with mean 00 and variance $1$ with law $\mathbb{P}_{X}$. Combining (5.2) and (5.3) shows that
$\mathbb{P}\otimes\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathcal{B})\geq\sum_{\omega\in\mathscr{S}_{\diamond}}\mathbb{P}\otimes\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathcal{B}\mid\omega)\cdot\phi_{p_{c},\Omega}^{\mathrm{f},0}(\omega)\geq(1-e^{-1}-0.1)\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathscr{S}_{\diamond}).$
The desired result comes from Theorem A.6 and Markov inequality. ∎
With Lemma 5.3, we are ready to prove Proposition 5.1.
Proof of Proposition 5.1.
Let $c_{1},c_{2}$ be defined in Lemma 5.3. For $h\in\mathcal{H}^{0}\cap\mathcal{H}_{\mathrm{cri}}(c_{1})$ (recall $\mathcal{H}^{0}$ from Definition 4.1) and $(h,\omega)\in\mathcal{B}$ (recall $\mathcal{B}$ from (5.1)), we have by Lemma 4.4
$\frac{\phi_{p_{c},\Omega}^{\mathrm{f},\epsilon h}}{\phi_{p_{c},\Omega}^{\mathrm{f},0}}=\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p_{c},\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p_{c},\Omega}}\cdot\prod_{\mathcal{C}\in\mathfrak{C}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T_{c}}\right)\leq e^{-1}\left(1+\sqrt{\epsilon N^{{\frac{7}{8}}}}\right)\leq\frac{1}{2}$
for sufficiently large $N$. Thus, we obtain that
$\|\phi_{p_{c},\Omega}^{\mathrm{f},\epsilon h}-\phi_{p_{c},\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq\frac{1}{2}\phi_{p_{c},\Omega}^{\mathrm{f},0}(\mathcal{B}\mid h)=\frac{1}{2}\phi_{p_{c},\Omega}^{\mathrm{f},0}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T_{c}}}\right)-\frac{\epsilon^{2}|\Omega|}{2T_{c}^{2}}\leq-1\right)\geq\frac{c_{1}}{2}.$
Combining Lemmas 4.3 and 5.3 and integrating over $\mathcal{H}^{0}\cap\mathcal{H}_{\mathrm{cri}}(c_{1})$, we obtain that
${\mathbb{E}}\|\phi_{p_{c},\Omega}^{\mathrm{f},\epsilon h}-\phi_{p_{c},\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq\frac{c_{1}}{2}\mathbb{P}\big{(}\mathcal{H}^{0}\cap\mathcal{H}_{\mathrm{cri}}(c_{1})\big{)}\geq\frac{c_{1}}{2}\left(c_{2}-c\exp\left(-c^{-1}\sqrt{\epsilon^{-1}N^{-\frac{7}{8}}}\right)\right)\geq\frac{c_{1}c_{2}}{3}$
for sufficiently large $N$. ∎
5.2 The high-temperature case
In this subsection, we consider the high-temperature case.
Proposition 5.4.
Fix $T>T_{c}(2)$ and thus $p<p_{c}$. Then there exists a constant $\theta_{0}=\theta_{0}(p)>0$ such that for any constant $\theta\in(0,\theta_{0})$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-\frac{1}{2}}\leq\epsilon\leq 2\theta N^{-\frac{1}{2}}$.
Remark 5.5.
One could easily remove the restriction of $\theta\in(0,\theta_{0})$ via a coarse graining method, see Section 6.1 for a detailed coarse graining framework (but with a different purpose).
Unlike the critical-temperature case, we do not have a good concentration bound for the ratio between partition functions of the FK-Ising model with and without disorder in the high-temperature case. So we focus on the fluctuation coming from all the clusters and show that this will lead to an anticoncentration of the measure. We start with some definitions.
Definition 5.6.
Let $\mathcal{F}\subset\{0,1\}^{\mathrm{E}(\Omega)}$ denote the set of edge configurations that $\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\leq N^{0.1}$ and $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\leq 2c_{1}N^{2}$ where $c_{1}$ comes from Lemma C.4. We use the notation $\mathcal{H}^{0}_{\mathrm{hig}}$ to denote the set of external field such that
$\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}\right)^{4}\Bigg{|}\mathcal{F}\Bigg{\rangle}\leq\exp\left(\epsilon^{-1}N^{-1/2}\right).$ (5.4)
Here and in the following, we write $\left\langle X\mid\mathcal{F}\right\rangle:=\frac{\left\langle X\cdot\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}{\left\langle\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}$ for any variable $X$ and event $\mathcal{F}$.
Definition 5.7.
For any constant $c>0$, we use the notation $\mathcal{H}_{\mathrm{hig}}^{1}(c)$ to denote the set of external field such that
$\phi_{p,\Omega}^{\mathrm{f},0}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\geq-c\epsilon^{4}N^{2}\right)\geq 1-N^{-1}.$ (5.5)
Further, for any constant $c>0$, let $\mathcal{H}_{\mathrm{hig}}^{2}(c)$ denote the collection of external field such that
$\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\geq c$
where $\mathsf{var}_{\phi}$ denotes the variance operator under $\phi_{p,\Omega}^{\mathrm{f},0}$ and $\mathsf{var}_{\phi}(\cdot|\mathcal{F})$ denotes the conditional variance operator with respect to $\mathcal{F}$.
Lemma 5.8.
There exists a constant $c>0$ depending only on $p$ such that for any disorder strength $\epsilon>0$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}\left(\mathcal{H}^{0}_{\mathrm{hig}}\right)\geq 1-c^{-1}\epsilon^{8}N^{4}\exp\left(-c{\epsilon^{-1}N^{-1/2}}\right).$
Lemma 5.9.
There exist constants $c_{1},c_{2}>0$ depending only on $p$ such that for any disorder strength $\epsilon>0$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}\left(\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\right){\geq 1-\frac{c_{2}}{N}}.$
Lemma 5.10.
There exists a constant $\theta_{0}<1$ such that the following holds. For any $\theta\leq\theta_{0}$, there exist constants $c_{3}=c_{3}(p),c_{4}=c_{4}(p)>0$ such that for any disorder strength $\theta N^{-1/2}\leq\epsilon\leq 2\theta N^{-1/2}$, domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}\left(\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})\right)\geq c_{4}.$
With Lemmas 5.8, 5.9 and 5.10, we are ready to prove Proposition 5.4.
Proof of Proposition 5.4.
To show the lower bound, we consider $h\in\mathcal{H}^{0}_{\mathrm{hig}}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ where $c_{1}$ and $c_{3}$ are the constants defined in Lemmas 5.9 and 5.10. Applying Lemmas 5.8, 5.9 and 5.10, it suffices to show that $\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq C$ for any $h\in\mathcal{H}_{\mathrm{hig}}^{0}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ and some constant $C>0$ independent of $h$. Let $X(\omega)=\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$. Applying (5.4), we have
$\langle X^{4}\mid\mathcal{F}\rangle\leq C_{1}.$ (5.6)
Since $h\in\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$, we have that
$\mathsf{var}_{\phi}(X\mid\mathcal{F})\geq c_{3}\theta^{4}.$ (5.7)
Combining (5.6), (5.7) and Corollary B.3, we get that there exists a constant $C_{2}>0$ such that for any real number $A$ there exists a set of configurations $\mathscr{S}_{1}=\mathscr{S}_{1}(A)$ with $\phi_{p,\Omega}^{\mathrm{f},0}(\mathscr{S}_{1}\mid\mathcal{F})\geq C_{2}$ such that for any $\omega\in\mathscr{S}_{1}(A)$, we have $|X(\omega)-A|\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}$. Applying Lemma C.4, we get that $\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{F})\geq\frac{1}{2}$ and thus $\phi_{p,\Omega}^{\mathrm{f},0}(\mathscr{S}_{1})\geq\frac{C_{2}}{2}$. Let $\mathscr{S}_{2}$ denote the collection of configurations such that $0\geq\sum_{\mathcal{C}\in\mathfrak{C}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|h_{\mathcal{C}}^{2}|}{2T^{2}}\geq-c_{1}\epsilon^{4}N^{2}$. Then for any $\omega\in\mathscr{S}_{2}$, we have
$\Bigg{|}\ln\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi_{p,\Omega}^{\mathrm{f},0}(\omega)}\right)\Bigg{|}=\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)+\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)\Bigg{|}$
$\geq-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}+\Bigg{|}X(\omega)+\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)+\frac{\epsilon^{2}|\Omega|}{2T^{2}}\Bigg{|}.$
Let $A=-\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$, then for any $\omega\in\mathscr{S}_{1}(A)\cap\mathscr{S}_{2}$, we have
$\Bigg{|}\ln\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi^{\mathrm{f},0}_{p,\Omega}(\omega)}\right)\Bigg{|}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-c_{1}\epsilon^{4}N^{2}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-16c_{1}\theta^{4}.$
Note that $f(x)\leq\frac{x^{2}}{2}$, then (5.5) shows that ${\phi_{p,\Omega}^{\mathrm{f},0}}\left(\mathscr{S}_{2}\right)\geq 1-N^{-1}$ and thus ${\phi_{p,\Omega}^{\mathrm{f},0}}\left(\mathscr{S}_{1}(A)\cap\mathscr{S}_{2}\right)\geq C_{3}>0$ for $N$ large enough. Hence we complete the proof of Proposition 5.4 by letting $\theta_{0}>0$ small enough. ∎
Now we turn to the proofs of Lemmas 5.8, 5.9 and 5.10.
Proof of Lemma 5.8.
For any configuration $\omega\in\mathcal{F}$, we obtain that
${\mathbb{E}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}\right)^{4}$
[…]
Summing over all configurations over $\mathcal{F}$ shows that
${\mathbb{E}}~{}\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}\right)^{4}\Bigg{|}\mathcal{F}\Bigg{\rangle}\leq C_{2}\epsilon^{8}N^{4},$
and the desired result comes from Markov inequality. ∎
Proof of Lemma 5.9.
Let $\mathcal{B}\subset\mathbb{R}^{\Omega}\otimes\{0,1\}^{\mathrm{E}(\Omega)}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\geq-C_{1}\epsilon^{4}N^{2}$ where $C_{1}>0$ is some constant to be determined. Since $f(x)\geq\frac{x^{2}-x^{4}}{2}$, we obtain that for any $\omega\in\mathcal{F}$,
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{B}^{c}\mid\omega)$ $\leq\mathbb{P}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}h_{\mathcal{C}}^{4}}{2T^{4}}>C_{1}\epsilon^{4}N^{2}\right)=\mathbb{P}_{X}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}X_{\mathcal{C}}^{4}}{2T^{4}}>C_{1}\epsilon^{4}N^{2}\right)$ (5.8)
where $\{X_{\mathcal{C}}\}_{\mathcal{C}\in\mathfrak{C}}$ are independent Gaussian variable with mean 00 and variance $1$. To compute the right-hand side of (5.8), we compute its first moment and variance. Since ${\mathbb{E}}_{X}X_{\mathcal{C}}^{4}=3$ and recall $\omega\in\mathcal{F}$, we get that
${\mathbb{E}}_{X}\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}X_{\mathcal{C}}^{4}}{2T^{4}}=\sum_{\mathcal{C}\in\mathfrak{C}}\frac{3\epsilon^{4}|\mathcal{C}|^{2}}{2T^{4}}\leq C_{2}\epsilon^{4}N^{2}.$ (5.9)
Moreover, we have
$\mathsf{var}_{X}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}X_{\mathcal{C}}^{4}}{2T^{4}}\right)=\sum_{\mathcal{C}\in\mathfrak{C}}{\mathbb{E}}_{X}~{}\frac{\epsilon^{8}|\mathcal{C}|^{4}(X_{\mathcal{C}}^{4}-3)^{2}}{4T^{8}}\leq C_{3}\epsilon^{8}N^{2}.$ (5.10)
Combining (5.9), (5.10) with (5.8) and letting $C_{1}>2C_{2}$, we get that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{B}^{c}\mid\omega)\leq\frac{C_{4}}{N^{2}}.$
Integrating over $\omega\in\mathcal{F}$ and applying Lemma C.4 shows that $\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{B}^{c})\leq C_{5}\exp\left(-C_{5}^{-1}N^{0.05}\right)+\frac{C_{4}}{N^{2}}.$ The desired result on $\mathbb{P}(\mathcal{H}_{\mathrm{hig}}^{1}(C_{1}))$ thus comes from the Markov inequality and letting $N$ big enough. ∎
Proof of Lemma 5.10.
Let $c_{1}>0$ be a constant to be defined later. By Lemma C.4, we get that
$\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{F})\geq 1-\exp(-CN^{0.05})$ (5.11)
for some constant $C>0.$ To lower-bound $\mathbb{P}(\mathcal{H}_{\mathrm{hig}}^{2}(c))$, we proceed by lower-bounding the first moment of $\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\mid\mathcal{F}\right)$ and also upper-bounding its second moment. We first expand the variance into
$\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)=\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\right)^{2}\Bigg{|}\mathcal{F}\Bigg{\rangle}-\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\right)\cdot\left(\sum_{\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}\frac{\epsilon^{2}h_{\tilde{\mathcal{C}}}^{2}}{2T^{2}}\right)\Bigg{|}\mathcal{F}\Bigg{\rangle}$
where ${\mathfrak{C}},\tilde{\mathfrak{C}}$ are the collections of all clusters of two independently sampled configurations according to $\phi_{p,\Omega}^{\mathrm{f},0}(\cdot\mid\mathcal{F})$. Thus, we compute
${\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)={\mathbb{E}}~{}\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\right)^{2}\Bigg{|}\mathcal{F}\Bigg{\rangle}-{\mathbb{E}}\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\right)\cdot\left(\sum_{\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}\frac{\epsilon^{2}h_{\tilde{\mathcal{C}}}^{2}}{2T^{2}}\right)\Bigg{|}\mathcal{F}\Bigg{\rangle}$
[…]
where the last equation comes from the fact that ${\mathbb{E}}h_{A}^{2}=|A|$ and ${\mathbb{E}}h_{A}^{2}h_{B}^{2}=|A|\cdot|B|+2|A\cap B|^{2}$. Combined with Cauchy inequality, it yields that
${\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)$ $\geq C_{1}\epsilon^{4}\left\langle\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\Big{|}\mathcal{F}\right\rangle$
$\geq C_{1}\epsilon^{4}\left\langle\left[\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\right]\cdot\mathbf{1}_{\mathcal{F}}\right\rangle$ (5.12)
Note that for $\omega\in\mathcal{F}^{c}$, we have the bound $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq(2N)^{4}.$ Combined with (5.12), it yields that
${\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\geq C_{1}\epsilon^{4}\left\langle\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\right\rangle-C_{1}\epsilon^{4}(2N)^{4}\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{F}^{c}).$ (5.13)
Plugging (5.11) and Lemma C.6 into (5.13), we get that
${\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\geq C_{2}\epsilon^{4}N^{2}.$ (5.14)
Next, we compute the upper bound of the second moment. We upper-bound the variance by
$\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\leq\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}\right)^{2}\Bigg{|}\mathcal{F}\Bigg{\rangle}.$
Applying Cauchy inequality, we obtain that
${\mathbb{E}}~{}\left(\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Big{|}\mathcal{F}\right)\right)^{2}\leq\Bigg{\langle}{\mathbb{E}}~{}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}\right)^{4}\Big{|}\mathcal{F}\Bigg{\rangle}$
[…]
Note that ${\mathbb{E}}~{}(h_{A}^{2}-|A|)^{4}=60|A|^{4}$ and ${\mathbb{E}}~{}(h_{A}^{2}-|A|)^{2}=2|A|^{2}$. Thus, we get that
${\mathbb{E}}~{}\left(\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\right)\Bigg{|}\mathcal{F}\right)^{2}$ $\leq C_{3}\epsilon^{8}\Bigg{\langle}\left(\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\right)^{2}\Bigg{|}\mathcal{F}\Bigg{\rangle}$ (5.15)
Recall the definition of $\mathcal{F}$ in Definition 5.6 , we get for any $\omega\in\mathcal{F}$ that $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\\ \leq cN^{2}.$ Combined with (5.15), it yields that
${\mathbb{E}}~{}\left(\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\right)^{2}\leq C_{4}\epsilon^{8}N^{4}.$ (5.16)
Combining (5.14), (5.16) and Paley-Zygmund inequality, we obtain that
$\mathbb{P}\left(\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\geq\frac{1}{2}{\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\right)\geq~{}\frac{\frac{1}{4}\left[{\mathbb{E}}~{}\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\right]^{2}}{{\mathbb{E}}~{}\left(\mathsf{var}_{\phi}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}\mathcal{F}\right)\right)^{2}}\geq\frac{C_{2}^{2}}{4C_{4}}.$
The desired comes from (5.14) and the fact that $\theta N^{-\frac{1}{2}}\leq\epsilon\leq 2\theta N^{-\frac{1}{2}}$. ∎
5.3 The low-temperature case
In this section, we consider the low-temperature case and fix $p>p_{c}$. The main goal of this subsection is to prove that the good external field in Definition 6.7 has a positive $\mathbb{P}$ probability not depending on its size. The strategy we use in the low-temperature regime is similar to that in the high-temperature regime. The difference is that in the high-temperature case, all the clusters are small such that none of them has a typical fluctuation, so they will contribute together and lead to the fluctuation. However, in the low-temperature case, the maximal cluster itself can lead to a big fluctuation. We start with some definitions as in Definitions 5.6 and 5.7, which was hinted in Definition 6.7. For the sake of convenience, we divide the definition of a good external field into the following three sets.
Definition 5.11.
We use the notation $\mathcal{H}_{\mathrm{low}}^{0}$ to denote the set of external field such that
$\Big{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\Big{\rangle}\leq\exp(\epsilon^{-1}N^{-1})~{}\text{and}~{}$ (5.17)
$\Big{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\Big{\rangle}\leq\exp(\epsilon^{-1}N^{-1}).$ (5.18)
Recall that $f(x)=\ln\cosh(x)$ and $\mathcal{C}_{*}$ is the boundary cluster. We use the notation $\mathcal{H}_{\mathrm{low}}^{1}$ to denote the set of external field such that
$\phi_{p,\Omega}^{\mathrm{w},0}\left(\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}(|\Omega|-|\mathcal{C}_{*}|)}{2T^{2}}~{}\Bigg{|}\leq N^{-0.1}\right)\geq 1-N^{-0.1}.$ (5.19)
For any constant $c>0$, let $\mathcal{H}_{\mathrm{low}}^{2}(c)$ denote the set of external field such that $\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\geq c$ where $\mathsf{var}_{\phi}$ denotes the variance operator under $\phi_{p,\Omega}^{\mathrm{w},0}$.
Lemma 5.12.
There exists a constant $c>0$ depending only on $p$ such that for any disorder strength $\epsilon>0$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}(\mathcal{H}^{0}_{\mathrm{low}})\geq 1-c\exp\left(-\epsilon^{-1}N^{-1}+c\epsilon^{2}N^{2}\right).$
Proof.
For any configuration $\omega$, let $X$ denote a standard Gaussian variable and we have
${\mathbb{E}}~{}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)={\mathbb{E}}_{X}~{}\exp\left(\frac{\epsilon\sqrt{|\mathcal{C}_{*}|}|X|}{T}\right)\leq\exp\left(C_{1}\epsilon^{2}|\mathcal{C}_{*}|\right)\mbox{ and }$
${\mathbb{E}}~{}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)={\mathbb{E}}_{X}~{}\cosh\left(\frac{\epsilon\sqrt{|\mathcal{C}|}X}{T}\right)\leq\exp(C_{1}\epsilon^{2}|\mathcal{C}|)$
for some absolute constant $C_{1}>0.$ Summing over all configurations shows that
${\mathbb{E}}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\Bigg{\rangle}$ $\leq{\mathbb{E}}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\Bigg{\rangle}$
$\leq\Big{\langle}\exp(C_{1}\epsilon^{2}|\mathcal{C}_{*}|)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\exp(C_{1}\epsilon^{2}|\mathcal{C}|)\Big{\rangle}\leq\exp(4C_{1}\epsilon^{2}N^{2}).$
The desired results come from an application of the Markov inequality. ∎
Lemma 5.13.
There exist constants $\theta_{0}>0$ and $c=c(p)>0$ such that the following holds. For any $0<\theta\leq\theta_{0}$, any disorder strength $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}(\mathcal{H}_{\mathrm{low}}^{1})\leq{1-c^{-1}\exp(-cN^{c})}.$
Proof.
We choose $\theta_{0}=1$. Let $\mathcal{B}^{+}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\leq N^{-0.1}$ and $\mathcal{B}^{-}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\geq-N^{-0.1}$. We first control the probability of $\mathcal{B}^{+}$ under $\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}$. For any configuration $\omega\in\mathscr{S}_{\mathrm{low},*}:=\{\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}|\mathcal{C}|\leq N^{0.1}\}$, we have
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\big{(}(\mathcal{B}^{+})^{c}\mid\omega\big{)}$ $=\mathbb{P}_{X}\left(\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\left(f\left(\frac{\epsilon\sqrt{|\mathcal{C}|}}{T}X_{\mathcal{C}}\right)-\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}\right)>N^{-0.1}\right)$
$\leq\mathbb{P}_{X}\left(\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)>N^{-0.1}\right)$
where the inequality comes from $f(x)\leq\frac{x^{2}}{2}$. Combined with Lemma B.5, it yields that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\left((\mathcal{B}^{+})^{c}\mid\omega\right)\leq c_{1}^{-1}\exp(-C_{1}N^{0.1}).$
Integrating over $\omega$ and applying Lemma C.1 shows that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\big{(}(\mathcal{B}^{+})^{c}\big{)}\leq C_{1}^{-1}\exp(-C_{1}N^{0.1})+\phi_{p,\Omega}^{\mathrm{w},0}(\mathscr{S}_{\mathrm{low},*}^{c})\leq C_{2}^{-1}\exp(-C_{2}N^{C_{2}}).$
Next we control $\mathcal{B}^{-}$. For any configuration $\omega\in\mathscr{S}_{\mathrm{low},*}$, we have
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\big{(}(\mathcal{B}^{-})^{c}\mid\omega\big{)}=\mathbb{P}_{X}\left(\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon\sqrt{|\mathcal{C}|}}{T}X_{\mathcal{C}}\right)-\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}<-N^{-0.1}\right)$
[…]
where the inequality comes from $f(x)\geq\frac{x^{2}-x^{4}}{2}$. Combined with Lemma B.5, it yields that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\big{(}(\mathcal{B}^{-})^{c}\mid\omega\big{)}\leq C_{3}^{-1}\exp(-C_{3}N^{0.1}).$
Integrating over $\omega$ and applying Lemma C.1 shows that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}\big{(}(\mathcal{B}^{-})^{c}\big{)}\leq C_{3}^{-1}\exp(-C_{3}N^{0.1})+\phi_{p,\Omega}^{\mathrm{w},0}(\mathscr{S}_{\mathrm{low},*}^{c})\leq C_{4}^{-1}\exp(-C_{4}N^{C_{2}}).$
The desired result on $\mathbb{P}(\mathcal{H}_{\mathrm{low}}^{1})$ thus comes from Markov inequality. ∎
Lemma 5.14.
For any $\theta>0$, there exist constants $c_{1}=c_{1}(p),c_{2}=c_{2}(p)>0$ such that for any disorder strength $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$, domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\mathbb{P}(\mathcal{H}_{\mathrm{low}}^{2}(c_{1}\theta^{2}))\geq c_{2}.$
Proof of Lemma 5.14.
Let $c>0$ be a constant to be defined later. To lower-bound $\mathbb{P}(\mathcal{H}_{\mathrm{low}}^{2}(c_{1}\theta))$, we proceed by lower-bounding the first moment of $\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}$ and also upper-bounding its second moment. We first expand the variance into
$\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}=\big{\langle}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}-\big{\langle}\epsilon h_{\mathcal{C}_{*}}\cdot\epsilon h_{\tilde{\mathcal{C}_{*}}}\big{\rangle}$
where $\mathcal{C}_{*},\tilde{\mathcal{C}}_{*}$ are the maximal clusters of two independently sampled configurations according to $\phi_{p,\Omega}^{\mathrm{w},0}$. Thus, we compute
${\mathbb{E}}~{}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}$ $=\big{\langle}{\mathbb{E}}~{}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}-\big{\langle}{\mathbb{E}}~{}\epsilon h_{\mathcal{C}_{*}}\cdot\epsilon h_{\tilde{\mathcal{C}_{*}}}\big{\rangle}=\big{\langle}\epsilon^{2}|\mathcal{C}_{*}|\big{\rangle}-\big{\langle}\epsilon^{2}|\mathcal{C}_{*}\cap\tilde{\mathcal{C}_{*}}|\big{\rangle}\geq C_{1}\epsilon^{2}N^{2}.$ (5.20)
where the last inequality comes from Lemma C.2. Next, we compute the upper bound of the second moment
${\mathbb{E}}~{}\Big{[}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{]}^{2}$ $\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}^{2}\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{4}h_{\mathcal{C}_{*}}^{4}\big{\rangle}=12\big{\langle}\epsilon^{4}|\mathcal{C}_{*}|^{2}\big{\rangle}\leq C_{2}\epsilon^{4}N^{4}.$ (5.21)
Combining (5.20), (5.21) and Paley-Zygmund inequality, we obtain that
$\mathbb{P}\Big{(}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\geq\frac{1}{2}{\mathbb{E}}~{}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{)}\geq\frac{\frac{1}{4}\Big{[}{\mathbb{E}}~{}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{]}^{2}}{{\mathbb{E}}~{}\Big{[}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{]}^{2}}\geq\frac{C_{1}^{2}}{4C_{2}}.$
The desired comes since $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$. ∎
Combining Lemmas 5.12, 5.14 and B.1, a similar version of Propositions 5.1 and 5.4 could be proved, and the proof is the same as that in Section 6.2. Thus, we just state the Proposition here and omit further details:
Proposition 5.15.
Fix $T<T_{c}(2)$ and thus that $p>p_{c}$. For any $\theta>0$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{w},\epsilon h}-\phi_{p,\Omega}^{\mathrm{w},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$.
5.4 Criticality of $\alpha(T)$
In this subsection, we will show that in order to make ${\mathbb{E}}\|\phi_{p,\Omega}^{\gamma,0}-\phi_{p,\Omega}^{\gamma,\epsilon h}\|_{\mathrm{TV}}$ converging to $1$, a necessary condition is that $\epsilon\gg N^{-\alpha(T)}$. More precisely, we will prove:
Proposition 5.16.
Fix $d=2$ and $T>0$. For any constant $\theta>0$, there exists a constant $c=c(\theta)>0$ such that ${\mathbb{E}}\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{\gamma,\epsilon h}\|_{\mathrm{TV}}\leq 1-c$ for any disorder strength $\epsilon\leq\theta N^{-\alpha(T)}$ and boundary condition $\gamma$.
We remark here that Proposition 5.16 is not a straightforward consequence of Section 4, because in Section 4, we focus on the region $\epsilon\ll N^{-\alpha(T)}$. Nevertheless, the proof idea of Proposition 5.16 is similar to that in Section 4. In order to overcome the difference that $\epsilon N^{\alpha(T)}$ might be big, we need to modify the definitions in Section 4.
Definition 5.17.
Let $\gamma\in\{-1,1\}^{\partial_{int}\Lambda_{N}}$ be a boundary condition on $\partial_{int}\Lambda_{N}$. Let $\mathcal{H}^{i}_{\diamond}(c)\subset{\mathbb{R}}^{\Lambda_{N}}~{}(i=1,2,3)$ denote the collection of the external field $h$ such that the followings hold
  1. 1.
    $c^{-1}\leq\frac{\mathcal{Z}^{\gamma,\epsilon h}_{\phi,p,\Omega}}{\mathcal{Z}^{\gamma,0}_{\phi,p,\Omega}}\cdot\frac{1}{\exp(\epsilon^{2}|\Omega|/2T^{2})}\leq c.$ (5.22)
  2. 2.
    $\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\leq c\right)\geq 1-c^{-1}\,.$ (5.23)
  3. 3.
    $\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\geq-c\right)\geq 1-c^{-1}\,.$ (5.24)
Following the idea in Section 4.2, it is easy to prove that there exists a constant $c>0$ depending on $\theta$ and $T$ such that for $i=2,3$ we have
$\mathbb{P}\big{(}\mathcal{H}^{i}_{\diamond}(c)\big{)}\geq 0.9\,.$
For $\mathcal{H}^{1}_{\diamond}$, it is a little more complicated since (4.4) no longer provides a lower bound on the ratio of the partition function when $\epsilon N^{\beta(T)}$ is bigger than $1$. The idea is that although we cannot directly control the ratio of the partition function, we can divide $\Lambda_{N}$ into small areas and show that with high $\mathbb{P}$-probability adding an external field on each small area perturbs the ratio by at most a factor of $2$. We summarize as the following Lemma.
Lemma 5.18.
[15, Lemma 3.4] Let $\Omega$ be an $M$-box such that $\Omega\subset\Lambda_{N}$, $\gamma\in\{-1,1\}^{\partial_{int}\Lambda_{N}}$ be the boundary condition and $g\in\mathbb{R}^{\Lambda_{N}\setminus\Omega}$ be the external field on $\Lambda_{N}\setminus\Omega$. Let $\mathcal{H}^{\gamma,g}_{\emptyset}\subset{\mathbb{R}}^{\Omega}$ denote the collection of the external field $h$ such that
$\Bigg{|}\Bigg{\langle}~{}\prod_{v\in\Omega}\left[1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\right]~{}\Bigg{\rangle}^{\gamma,g}_{p,\Lambda_{N}}-1\Bigg{|}\leq\sqrt{\epsilon M^{\beta(T)}}.$ (5.25)
There exist constants $c_{1},c_{2}>0$ depending only on $T$ such that
$\mathbb{P}(\mathcal{H}^{\gamma,g}_{\emptyset})\geq 1-c_{1}\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}M^{-\beta(T)}}\right)$
for all disorder strength $\epsilon\leq c_{2}M^{-\beta(T)}$.
For any disorder strength such that $\epsilon\leq\theta N^{-\alpha(T)}$, let $M=\lfloor C_{1}\epsilon^{-\frac{1}{\alpha(T)}}\rfloor$ with $C_{1}>0$ small enough. Let $\{B_{i}\}_{i\in I}$ be a partition of $\Lambda_{N}$ by $M$-boxes, then we can conduct an induction by Lemma 5.18 and obtain that
$\mathbb{P}(\mathcal{H}_{\diamond}^{1}(2^{|I|}))\geq 1-c_{1}|I|\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}M^{-\beta(T)}}\right).$
Note that $|I|=\frac{N^{2}}{M^{2}}\leq M^{-2}(\theta\epsilon^{-1})^{\frac{2}{\alpha(T)}}$, letting $C_{1}$ small enough shows that $\mathbb{P}\left({\mathcal{H}_{\diamond}^{1}(2^{|I|})}\right)\geq 0.9$. The rest of the proof is similar to that in Section 4 and thus we omit further details.
6 Singularity in strong-disorder regime
This section is devoted to the proof of singularity results in Theorem 1.3. It combines a coarse-graining argument with some near-critical results that have be shown in Section 5. The proof idea is quite similar to Section 3.2, however, careful analysis of the random cluster configuration is involved since the RFFKIM lacks the nice domain Markov property (DMP) possessed by the RFIM.
6.1 Critical-temperature and high-temperature cases
Critical-temperature and high-temperature regimes can be analyzed jointly. The core intuition is that in both cases, random clusters remain bounded in size – equivalently, dual paths are abundant.
More precisely, let us consider the boxes inside $\Lambda_{N}$ that are surrounded by closed edges. Let $M$ be an integer with order $O(\epsilon^{-\frac{1}{\alpha(T)}})$ whose value will be defined later and without loss of generality, we assume $N$ is divisible by $2M$. Let $B_{i}~{}(i=1,\cdots,n)$ be a partition of $\Lambda_{N}$ by $2M$-boxes and let $u_{i}$ denote the center of $B_{i}$ (i.e. $B_{i}=\Lambda_{2M}(u_{i})$ and $n=\Big{(}\frac{N}{2M}\Big{)}^{2}$).
Definition 6.1.
For any configuration $\omega\in\{0,1\}^{\rm{E}}$, we say $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n})$ is the outmost close region of $\omega$ if
  1. (i)
    $\Omega_{i}\subset B_{i}$ and either $\Lambda_{M}(u_{i})\subset\Omega_{i}$ or $\Omega_{i}=\emptyset$ $(i=1,\cdots,n)$.
  2. (ii)
    $\forall e=\{x,y\}\in\rm{E}$, $x\in\Omega_{i},y\notin\Omega_{i}\,$, $e$ is close in $\omega$.
  3. (iii)
    There does not exists a subset $\Lambda_{M}(u_{i})\subset\Omega_{i}^{\prime}\subset\Lambda_{2M}(u_{i})$ satisfying (ii) such that $\Omega_{i}\subsetneqq\Omega_{i}^{\prime}$.
See Figure 1 for illustration. We use the notation $\mathcal{O}=\mathrm{Out}(\omega)$ to denote the event that $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n})$ is the outmost close region of $\omega$. Let $\eta(\mathcal{O})$ denote the number of nonempty regions $\Omega_{i}$, i.e. $\eta(\mathcal{O})=\sum_{i=1}^{n}\mathbf{1}_{\Omega_{i}\neq\emptyset}$ and $\Omega(\mathcal{O})=\bigcup_{i=1}^{n}\Omega_{i}$ denote the union of the regions.
Figure 1: An illustration of the outmost close region. Dotted lines: the dual circuits.
For any $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n})$, let $\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})=\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\omega)$ be the marginal distribution on the outmost close region. We emphasize that the event $\mathcal{O}=\mathrm{Out}(\omega)$ is measurable under $\omega_{|E(\Lambda_{N})\setminus E(\Omega(\mathcal{O}))}$. We further define $\Phi_{p,\mathcal{O}}^{h}$ to be the product measure of $\phi^{\mathrm{f},h}_{p,\Omega_{i}}~{}(i=1,2,\cdots,n)$. More precisely, let $\omega^{\mathcal{O}}\in\{0,1\}^{E(\Omega(\mathcal{O}))}$ be an edge configuration and $\omega_{i}^{\mathcal{O}}=\omega^{\mathcal{O}}_{|\Omega_{i}},$ then we define $\Phi_{p,\mathcal{O}}^{h}(\omega^{\mathcal{O}})=\prod_{i=1}^{n}\phi_{p,\Omega_{i}}^{\mathrm{f},h}(\omega_{i}^{\mathcal{O}}).$ By applying DMP, we obtain that for any configuration $\omega$ and $\mathcal{O}=\mathrm{Out}(\omega)$,
$\phi_{p,\Lambda_{N}}^{\gamma,0}(\omega)$ $=\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{O})\cdot\Phi_{p,\mathcal{O}}^{0}(\omega^{\mathcal{O}}),~{}~{}~{}~{}\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\omega)=\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})\cdot\Phi_{p,\mathcal{O}}^{\epsilon h}(\omega^{\mathcal{O}}).$ (6.1)
We first claim that in order to show $\|\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}-\phi_{p,\Lambda_{N}}^{\gamma,0}\|_{\mathrm{TV}}$ is close to $1$, it suffices to show that $\|\Phi_{p,\mathcal{O}}^{\epsilon h}-\Phi_{p,\mathcal{O}}^{0}\|_{\mathrm{TV}}$ is close to $1$ with high probability under the measure $\phi_{p,\Lambda_{N}}^{\gamma,0}$ for overwhelming proportion of external field $\epsilon h$. To see this, for any fixed external field $\epsilon h,$ let $\mathscr{O}=\mathscr{O}(\epsilon h)$ denote the set of $\mathcal{O}$ such that
$\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\geq 1-c_{1}\exp(-c_{1}^{-1}n).$ (6.2)
Similar to (3.5) and (3.6), we apply (6.1) to obtain that
$1-\|\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}-\phi_{p,\Lambda_{N}}^{\gamma,0}\|_{\mathrm{TV}}\,{=}\sum_{\mathcal{O}}\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\min\left\{\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{O})\cdot\Phi_{p,\mathcal{O}}^{0}(\omega),\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})\cdot\Phi_{p,\mathcal{O}}^{\epsilon h}(\omega)\right\}$
$\leq$ $\sum_{\mathcal{O}\in\mathscr{O}}\big{(}\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{O})+\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})\big{)}\cdot\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\min\left\{\Phi_{p,\mathcal{O}}^{0}(\omega),\Phi_{p,\mathcal{O}}^{\epsilon h}(\omega)\right\}+$
$\sum_{\mathcal{O}\in\mathscr{O}^{c}}\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{O})\cdot\Phi_{p,\mathcal{O}}^{0}(\omega)$
[…]
This naturally leads to the following definition:
Definition 6.2.
We use the definition $\mathcal{H}_{\diamond}(c)$ to denote the set of external field such that
$\phi_{p,\Lambda_{N}}^{\gamma,0}\Big{(}\big{\{}\mathcal{O}:\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\geq 1-c\exp(-c^{-1}n)\big{\}}\Big{)}\geq 1-c\exp(-c^{-1}n).$ (6.4)
Lemma 6.3.
Fix temperature $T\geq T_{c}$ and $\theta>0$, there exist constants $c_{1},c_{2}>0$ only relies on $T,\theta$ such that for any disorder strength $\theta M^{-\alpha(T)}\leq\epsilon\leq 2\theta M^{-\alpha(T)}$, we have
$\mathbb{P}\big{(}\mathcal{H}_{\diamond}(c_{1})\big{)}\geq 1-c_{2}\exp\left(-c_{2}^{-1}\left(\frac{N}{2M}\right)^{2}\right).$
Proof.
Recall that $\{B_{i}\}_{1\leq i\leq n}$ is a partition of $\Lambda_{N}$ by $2M$-boxes. Applying the RSW theory (see [22] and also e.g. [21, 36, 29] for other remarkable progress on the RSW theory), there exists a constant $C_{1}>0$ such that $\phi^{\xi,0}_{p,B_{i}}(\Omega_{i}\neq\emptyset)\geq C_{1}$ where $\Omega_{i}$ is the region satisfying Definition 6.1 and $\xi$ is an arbitrary boundary condition on $\partial_{int}B_{i}$. Thus, we obtain by DMP and Lemma B.6 that
$\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\left\{\eta(\mathcal{O})\geq\frac{C_{1}n}{2}\right\}\right)\geq 1-C_{2}\exp({-}C_{2}^{-1}n).$ (6.5)
Let $\mathcal{A}$ denote the event that $1-\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\leq C_{3}\exp({-}C_{3}^{-1}n)$ with some $C_{3}>0$ to be determined later. Note that $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$ are product measures of $\phi^{\mathrm{f},0}_{p,\Omega_{i}}$ and $\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}$ and Propositions 5.1 and 5.4 show that ${\mathbb{E}}\|\phi^{\mathrm{f},0}_{p,\Omega_{i}}-\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}\|_{\mathrm{TV}}\geq C_{4}>0$. Thus, applying Corollary 3.4 gives that for any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq\frac{C_{1}n}{2}$
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}\mid\{\mathcal{O}=\mathrm{Out}(\omega)\})\geq 1-C_{5}\exp({-}C_{5}^{-1}n).$ (6.6)
Note that we also fix $C_{3}$ here to be the constant given by Corollary 3.4.
Integrating (6.6) over $\{\eta(\mathcal{O})\geq\frac{C_{1}n}{2}\}$ and combining with (6.5), we get that
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}^{c})\leq C_{6}\exp({-}C_{6}^{-1}n).$ (6.7)
Thus, the desired result comes from Markov inequality and letting $c_{1}>\max\{C_{3},C_{6}\}$. ∎
With Lemma 6.3, we are ready to prove the second criterion of Theorem 1.3.
Proof of Theorem 1.3 (high-temperature and critical-temperature) part two.
Let $M$ be an integer such that $\theta M^{-\alpha(T)}\leq\epsilon\leq 2\theta M^{-\alpha(T)}$ and we can simply choose $\theta=1$. We first consider $h\in\mathcal{H}_{\diamond}(c_{1})$ where $c_{1}$ was defined in Lemma 6.3. Combining (6.3) with (6.4) and (6.2), we get that
$1-\|\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}-\phi_{p,\Lambda_{N}}^{\gamma,0}\|_{\mathrm{TV}}$ $\leq\sum_{\mathcal{O}\in\mathscr{O}}\big{(}\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{O})+\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})\big{)}\cdot c_{1}\exp(-c_{1}^{-1}n)+c_{1}\exp(-c_{1}^{-1}n)$
$\leq 3c_{1}\exp(-c_{1}^{-1}n).$
The desired result comes from Lemma 6.3. ∎
6.2 The low-temperature case
Now we consider the low-temperature case. Let $M<N$ be a fixed integer with order $O(\epsilon^{-1})$ to be defined later. The main difference from Section 6.1 is that (6.5) does not hold since with high probability there does not exist a dual circuit in $\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i})$. Thus, we need to consider an open circuit instead of a dual circuit, which provides a wired boundary condition instead of a free boundary condition. Further, we emphasize that the second equality in (6.1) fails since the external field outside $\Omega_{i}$ also influences the measure. Nevertheless, the influence of outside external field is manageable if the open circuit is connected to a unique large cluster.
We start with the construction of the desired unique large cluster.
Definition 6.4.
Recall that $\{B_{i}=\Lambda_{2M}(u_{i}):i=1,2,\cdots,n\}$ is a partition of $\Lambda_{N}$ by $2M$-boxes. Without loss of generality, we assume $M$ is an even number and $N$ is divisible by $2M$. Define $\Lambda:=\bigcup_{i=1}^{n}(\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i}))$ to be the union of the annuli in each $B_{i}$ and we point out that $\Lambda$ is connected.
An edge configuration $\omega$ is called well-connected if there exists a unique connected cluster in $\omega_{|\Lambda}$ such that its diameter is not less than $N/2.$ And we will call this cluster the main cluster in $\omega$ for convenience. Besides, we write $\mathcal{W}\subset\{0,1\}^{\mathbb{E}(\Lambda_{N})}$ for the set of well-connected configurations.
Remark 6.5.
We highlight that in the definition of a well-connected configuration $\omega$, the uniqueness of the main cluster is confined to $\omega_{\mid\Lambda}$. In other words, despite this restriction, there remains the possibility that certain small clusters within $\omega_{\mid\Lambda}$ may be connected to a large cluster in $\omega$ that differs from the main cluster. This clarification is solely intended to enhance readers’ understanding and does not impact the proof in any way.
Then we will explore the open circuits in each $B_{i}$ that is connected to the main cluster. The following definition comes from a similar idea as Definition 6.1.
Definition 6.6.
For any well-connected configuration $\omega\in\mathcal{W}$ and any external field $h$, we say $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})$ is the outmost good open region of $\omega$ if
  1. (i)
    For all $i=1,2,\cdots,n,$ either $\Omega_{i}=\emptyset$ or $\Lambda_{M/2}(u_{i})\subset\Omega_{i}\subset\Lambda_{M}(u_{i})$ such that the circuit boundary of $\Omega_{i}$ is open and it is connected to the main cluster of $\omega.$
  2. (ii)
    If $\Omega_{i}\neq\emptyset$, then $\epsilon h$ is $c$-good with respect to $\Omega_{i}$ where the $c$-good external field will be defined in the Definition 6.7 below.
  3. (iii)
    There does not exists a subset $\Lambda_{M/2}(u_{i})\subset\Omega^{\prime}_{i}\subset\Lambda_{M}(u_{i})$ satisfying (i) and (ii) such that $\Omega_{i}\subsetneqq\Omega^{\prime}_{i}$.
  4. (iv)
    Let $\Gamma=\cup_{i=1}^{n}(B_{i}\setminus\Omega_{i})$, then $\tilde{\omega}=\omega_{|\Gamma}$.
See Figure 2 for illustration. Let $\mathcal{C}$ be the main cluster of $\tilde{\omega}$ in $\Lambda$. Let $\mathcal{C}^{\prime}$ be the cluster that contains $\mathcal{C}$ of $\tilde{\omega}$ in $\Gamma$. Note that $\mathcal{C}\subset\mathcal{C}^{\prime}$. We slightly abuse the notation and call $\mathcal{C}^{\prime}$ the main cluster in $\tilde{\omega}$.
Figure 2: An illustration of the outmost good open region. The dotted squares are $\Lambda_{M}(u_{i})$ where $\Lambda_{2M}(u_{i})=B_{i}$. We do not draw the box $\Lambda_{M/2}(u_{i})$, but each red and orange region contains a box of such size. The red color denotes that $h$ is $c$-good with respect to the region, and the orange color denotes that $h$ is not $c$-good with respect to the region.
We use the notation $\mathcal{O}=\mathrm{Out}(\omega,h)$ to denote that $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})$ is the outmost good open configuration of $\omega$. Note that for $\mathcal{O}=\mathrm{Out}(\omega,h)$, $\omega$ is well-connected if and only if $\tilde{\omega}$ is well-connected. Thus, we get that $\{\omega:\mathrm{Out}(\omega,h)=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})\}=\tilde{\omega}\otimes\{0,1\}^{E(\cup_{i=1}^{n}\Omega_{i})}$. Let $\eta(\mathcal{O})$ denote the number of nonempty regions $\Omega_{i}$. We define $\Phi_{p,\mathcal{O}}^{\epsilon h}$ to be FK measure $\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\cdot\mid\tilde{\omega})$ conditioned on $\omega_{|\Gamma}=\tilde{\omega}$. By calculations in (6.3), it suffices to prove a lower bound for $\|\Phi^{\epsilon h}_{p,\mathcal{O}}-\Phi^{0}_{p,\mathcal{O}}\|_{\mathrm{TV}}$ as in (6.2). Unlike Section 6.1, $\Phi^{\epsilon h}_{p,\mathcal{O}}$ is not a product measure of measures supported on $\Omega_{i}$ because the external field outside $\Omega_{i}$ will influence the measure on $\Omega_{i}$. Therefore, we introduce the following definition of good external field.
Definition 6.7.
For any region $\Lambda_{M/2}\subset\Omega\subset\Lambda_{M}$, consider the FK-Ising measure with wired boundary condition on $\Omega$. Let $\mathcal{C}_{*}$ be the boundary cluster, we say the external field $\epsilon h$ is $c$-good with respect to $\Omega$ if the following holds.
  1. (i)
    $\Big{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\Big{\rangle}_{p,\Omega}^{\mathrm{w},0}\leq\exp(\epsilon^{-1}M^{-1}).$
  2. (ii)
    $\Big{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\Big{\rangle}_{p,\Omega}^{\mathrm{w},0}\leq\exp(\epsilon^{-1}M^{-1}).$
  3. (iii)
    $\phi_{p,\Omega}^{\mathrm{w},0}\left(\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\leq M^{-0.1}\right)\geq 1-M^{-0.1}.$
  4. (iv)
    $\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\geq c$ where $\mathsf{var}_{\phi}$ denotes the variance operator under $\phi_{p,\Omega}^{\mathrm{w},0}$.
The following lemma provides a useful anticoncentration fact about the good external field with the proof postponed to the end of this section.
Lemma 6.8.
There exists a constant $\theta_{0}>0$ such that for any $0<\theta<\theta_{0}$ and $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$, we have the following. For any region $\Lambda_{M}\subset\Omega\subset\Lambda_{2M}$ and any $c$-good external field $\epsilon h$ with respect to $\Omega$, let $X^{\pm}=\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\pm\frac{\epsilon h_{\mathcal{C}_{*}}}{T}$. Then we have
$\langle X^{\pm}\rangle^{\mathrm{w},0}_{p,\Omega}\leq\ln\big{(}\langle\exp(X^{\pm})\rangle^{\mathrm{w},0}_{p,\Omega}\big{)}-c_{1}$
for some $c_{1}>0$ depending only on $\theta$ and $c$.
Then we claim that ${\eta(\mathcal{O})}$ is large for high $\mathbb{P}$-probability and also postpone the proof to the end of this section.
Definition 6.9.
We use the definition $\mathcal{H}_{\Box}(c)$ to denote the set of external field such that
$\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\left\{\mathcal{O}=\mathrm{Out}(\omega,h):{\eta(\mathcal{O})}\geq\frac{cN^{2}}{M^{2}}\right\}\right)\geq 1-c^{-1}\exp\left(-\frac{cN}{M}\right).$ (6.8)
Lemma 6.10.
Fix $T<T_{C}$ and $\theta<\theta_{0}$ where $\theta_{0}$ is defined in Lemma 6.8. Then there exist constants $c_{1},c_{2}>0$ such that for any disorder strength $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$, we have $\mathbb{P}\big{(}\mathcal{H}_{\Box}(c_{1})\big{)}\geq 1-c_{2}\exp\big{(}-c_{2}^{-1}\frac{N}{M}\big{)}.$
Now we are ready to give a similar bound as (6.2), as incorporated in the following Lemma 6.11.
Lemma 6.11.
For any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq c_{1}n$, we have
$\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\geq 1-c_{2}\exp(-c_{2}^{-1}n)$
where $c_{2}>0$ does not depend on $\mathcal{O}$ or $\epsilon h$.
Proof.
Fix any external field $\epsilon h$ and $\mathcal{O},$ let $I$ denote the set of indices such that $\Omega_{i}\neq\emptyset$ (thus $\epsilon h$ is good with respect to $\Omega_{i}$). For any $\mathrm{Out}(\omega,h)=\mathcal{O}$ and $i\in I$, let $\mathfrak{C}^{i}$ be the collection of clusters in $\omega_{\mid\Omega_{i}}$ and let $\mathcal{C}_{*}^{i}$ be the boundary cluster in $\Omega_{i}$. Let
$X_{i}^{\pm}(\omega)=\sum_{\mathcal{C}\in\mathfrak{C}^{i}\setminus\{\mathcal{C}_{*}^{i}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\pm\frac{\epsilon h_{\mathcal{C}_{*}^{i}}}{T}.$
Recall Definitions 6.4 and 6.6, we get that $\mathcal{C}_{*}^{i}$ are connected by the main cluster of $\omega$. Then we can compute the Radon-Nikodym derivative between $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$
$\frac{\Phi^{\epsilon h}_{p,\mathcal{O}}(\omega)}{\Phi^{0}_{p,\mathcal{O}}(\omega)}$ $=\frac{\prod_{i\in I}\prod_{\mathcal{C}\in\mathfrak{C}^{i}\setminus\{\mathcal{C}_{*}^{i}\}}\cosh(\frac{\epsilon h_{\mathcal{C}}}{T})\cdot\cosh\left(\sum_{i\in I}\frac{\epsilon h_{\mathcal{C}_{*}^{i}}}{T}+a\right)}{\Big{\langle}\prod_{i\in I}\prod_{\mathcal{C}\in\mathfrak{C}^{i}\setminus\{\mathcal{C}_{*}^{i}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\cdot\cosh\left(\sum_{i\in I}\frac{\epsilon h_{\mathcal{C}_{*}^{i}}}{T}+a\right)\Big{\rangle}^{0}_{p,\mathcal{O}}}$
$=\frac{\exp\left(\sum_{i\in I}X_{i}^{+}+a\right)+\exp\left(\sum_{i\in I}X_{i}^{-}-a\right)}{\Big{\langle}\exp\left(\sum_{i\in I}X_{i}^{+}+a\right)+\exp\left(\sum_{i\in I}X_{i}^{-}-a\right)\Big{\rangle}^{0}_{p,\mathcal{O}}}$
$=\frac{\exp\left(\sum_{i\in I}X_{i}^{+}+a\right)+\exp\left(\sum_{i\in I}X_{i}^{-}-a\right)}{\exp(a)\cdot\prod_{i\in I}\big{\langle}\exp(X_{i}^{+})\big{\rangle}^{\mathrm{w},0}_{p,\Omega_{i}}+\exp(-a)\cdot\prod_{i\in I}\big{\langle}\exp(X_{i}^{-})\big{\rangle}^{\mathrm{w},0}_{p,\Omega_{i}}}$ (6.9)
where $a$ is the sum of the external field on the main cluster of $\tilde{\omega}$. Recalling Definition 6.9 and combining with Lemma 6.8, we get that
$\langle X_{i}^{\pm}\rangle^{0}_{p,\Omega_{i}}\leq\ln\left(\langle\exp(X_{i}^{\pm})\rangle^{0}_{p,\Omega_{i}}\right)-C_{1}$
for some $C_{1}>0$. Let $\mathscr{S}^{\pm}$ denote the collection of configurations such that
$\big{|}\sum_{i\in I}X_{i}^{\pm}-\langle X_{i}^{\pm}\rangle^{0}_{p,\Omega_{i}}\big{|}\geq\frac{C_{1}n}{2}.$
Noticing that $\Phi^{0}_{p,\mathcal{O}}$ is the product measure of $\phi^{\mathrm{w},0}_{p,\Omega_{i}}$, by Lemma B.6 and $|I|\geq c_{1}n$ we obtain that
$\Phi^{0}_{p,\mathcal{O}}\Big{(}\mathscr{S}^{+}\cup\mathscr{S}^{-}\Big{)}\leq C_{2}^{-1}\exp(-C_{2}n).$ (6.10)
Further for any $\omega\in(\mathscr{S}^{+})^{c}\cap(\mathscr{S}^{-})^{c}$, we have that
$\exp\left(\sum_{i\in I}X_{i}^{\pm}\pm a\right)$ $\leq\exp\left(\sum_{i\in I}\left\langle X_{i}^{\pm}\right\rangle^{\mathrm{w},0}_{p,\Omega_{i}}\pm a+\frac{C_{1}n}{2}\right)$
$\leq\prod_{i\in I}\left\langle\exp(X_{i}^{\pm})\right\rangle^{\mathrm{w},0}_{p,\Omega_{i}}\cdot\exp\left(\pm a-\frac{C_{1}n}{2}\right).$ (6.11)
Combining (6.11) with (6.9) shows that for any $\omega\in(\mathscr{S}^{+})^{c}\cap(\mathscr{S}^{-})^{c}$
$\frac{\Phi^{\epsilon h}_{p,\mathcal{O}}(\omega)}{\Phi^{0}_{p,\mathcal{O}}(\omega)}\leq\exp\left(-\frac{C_{1}n}{2}\right).$ (6.12)
The desired result comes from combining (6.12) with (6.10). ∎
With Lemma 6.10 and Lemma 6.11 in place of Lemma 6.3, the proof of the singularity in the low-temperature regime is the same to that of high-temperature and critical-temperature regimes. Thus, we omit the details here.
Finally, we come back to the proof of Lemmas 6.8 and 6.10.
Proof of Lemma 6.8.
Combining (i) and (iv) in Definition 6.7 with Lemma B.1, we get that there exist two collections of configurations $\mathscr{S}_{1}$ and $\mathscr{S}_{2}$, such that $\phi_{p,\Omega}^{\mathrm{w},0}(\mathscr{S}_{i})\geq C_{1}$ and for any $\omega_{1}\in\mathscr{S}_{1},\omega_{2}\in\mathscr{S}_{2}$ we have $|\frac{\epsilon h_{\mathcal{C}_{*}(\omega_{1})}}{T}-\frac{\epsilon h_{\mathcal{C}_{*}(\omega_{2})}}{T}|\geq C_{2}$ where $C_{1},C_{2}>0$ are constants depending only on $\theta$ and $c$. Let $\mathscr{S}_{*}$ denote the collection of configurations such that
$\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\leq M^{-0.1}\text{ and }\big{|}~{}|\mathcal{C}_{*}|-\langle|\mathcal{C}_{*}|\rangle~{}\big{|}\leq M^{1.5}.$
Then letting $M$ big enough and combining (iii) in Definition 6.7 with Lemma C.3 shows that $\phi_{p,\Omega}^{\mathrm{w},0}(\mathscr{S}_{i}\cap\mathscr{S}_{*})\geq C_{3}$. Thus, for any $\omega_{1}\in\mathscr{S}_{1},\omega_{2}\in\mathscr{S}_{2}$ we have
$\big{|}X^{+}(\omega_{1})-X^{+}(\omega_{2})\big{|}\geq\Big{|}\frac{\epsilon h_{\mathcal{C}_{*}(\omega_{1})}}{T}-\frac{\epsilon h_{\mathcal{C}_{*}(\omega_{2})}}{T}\Big{|}-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}(\omega_{1})\setminus\{\mathcal{C}_{*}(\omega_{1})\}}f\left(\frac{\epsilon h_{\mathcal{C}(\omega_{1})}}{T}\right)-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}(\omega_{1})|}{2T^{2}}\Bigg{|}$
$\hskip 50.00008pt-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}(\omega_{2})\setminus\{\mathcal{C}_{*}(\omega_{2})\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}(\omega_{2})|}{2T^{2}}\Bigg{|}-\Bigg{|}\frac{\epsilon^{2}(\big{|}|\mathcal{C}_{*}(\omega_{1})|-|\mathcal{C}_{*}(\omega_{2})|\big{|})}{2T^{2}}\Bigg{|}\geq\frac{C_{2}}{2}.$
Therefore, we get that $\mathsf{var}_{\phi}(X^{+})\geq C_{4}$ for some $C_{4}>0$ depending only on $\theta$ and $c$. Combined with (ii) in Definition 6.7 and Lemma B.4, it yields that
$\langle X^{+}\rangle^{0}_{p,\Omega}\leq\ln\big{(}\langle\exp(X^{+})\rangle^{0}_{p,\Omega}\big{)}-C_{5}.$
By symmetry, the same result holds for $X^{-}$. ∎
Proof of Lemma 6.10.
We first prove that with high $\phi^{\gamma,0}_{p,\Lambda_{N}}$ probability, an edge configuration $\omega$ is well-connected. It is a typical result for the low-temperature FK-Ising model thus we just sketch the proof here.
For a box $B_{i}=\Lambda_{2M}(u_{i}),$ let $\mathcal{G}(u)$ denote the event that there exist two open circuits in the annuli $\Lambda_{3M}(u)\setminus\Lambda_{2M}(u)$ and $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$ respectively, and these two circuits are connected in $\omega{{}_{\mid\Lambda_{3M}(u)\setminus\Lambda_{M}(u)}}.$ Then by the RSW theory (see [22] for example), there exists a constant $C_{1}>0$ relying only on $p$ such that for any boundary condition $\gamma$ on $\partial_{int}\Lambda_{3M}(u_{i}),$ $\phi^{\gamma,0}_{p,\Lambda_{3M}(u_{i})}(\mathcal{G}(u))\geq 1-\exp(-C_{1}M).$ Therefore, we define $\mathcal{Q}(\omega)=\left(Q_{x}=\mathbf{1}_{\mathcal{G}(2Mx)}:x\in\Lambda_{\frac{N}{2M}}\right)$ to be a random vertex configuration on the box $\Lambda_{\frac{N}{2M}},$ and then by [32], we get that $\mathcal{Q}$ dominates a Bernoulli site percolation $\mathbb{P}_{q},$ where $q$ can be arbitrary close to 1 as $M$ goes to infinity. We define $\mathcal{A}$ to be the event that the following holds:
(1) There exists a unique cluster $C_{\mathcal{Q}}$ in $\mathcal{Q}$ with diameter at least $\frac{N}{4M}$ and volume at least $\frac{C_{2}N^{2}}{M^{2}}$ in $\mathcal{Q}(\omega)$.
(2) For any connected subset $S$ in $\Lambda_{\frac{N}{2M}}$ with diameter at least $\frac{N}{100M}$, we have $S\cap C_{\mathcal{Q}}\neq\emptyset.$
Next we will prove for any $\omega\in\mathcal{A}$, we have $\omega$ is well-connected. It is easy to see that $C_{\mathcal{Q}}$ induces a cluster $\mathcal{C}$ with diameter at least $N/2$ in $\Lambda=\bigcup_{i=1}^{n}(\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i}))$. So it suffices to prove that this is the unique cluster in $\Lambda$ with diameter at least $N/2$. For any cluster $\mathcal{C}^{\prime}$ in $\Lambda$ with diameter at least $N/2$, let $S=\{x\in\Lambda_{\frac{N}{2M}}:\Lambda_{M}(2Mx){\cap\mathcal{C}^{\prime}\neq\emptyset}\}$. By (2), we get that there exists $x\in S\cap C_{\mathcal{Q}}$. Since $Q_{x}=1$ and $x\in{S}$, $\mathcal{C}^{\prime}$ must intersect with the circuit in $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$. Thus we get $\mathcal{C}^{\prime}$ intersects with $\mathcal{C}$ and $\omega$ is well-connected. By standard percolation result, under the measure $\mathbb{P}_{p},$ $\mathcal{A}$ has probability $1-\exp(-C_{3}\frac{N}{M}).$ Thus, we get $\phi^{\gamma,0}_{p,\Lambda_{N}}(\omega\text{ is well-connected})\geq 1-\exp\left(-C_{4}\frac{N}{M}\right)$ by noticing that a connected component in $\mathcal{Q}$ induces an open cluster in $\omega.$
Figure 3: Illustration of the event $\mathcal{A}$. The blue circuit lies in $\Lambda_{2M}(x)\setminus\Lambda_{M}(x)$, the red circuit in $\Lambda_{3M}(x)\setminus\Lambda_{2M}(x)$, and the green path(s) connect these two circuits. Neighbouring red circuits intersect with each other.
We choose $\theta=1$ and $M$ such that $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$. Furthermore, we choose $c$ in Definition 6.7 to be the first constant in Lemma 5.14. Then by the results in Section 5.3 (see Lemmas 5.12, 5.13 and 5.14), we get that the external field on the region $\Omega$ is good with a positive probability. Hence by Lemma B.6, we get that
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}\big{(}\eta(\mathcal{O})\geq C_{5}n\big{)}\geq 1-C_{5}^{-1}\exp\left(-C_{5}\frac{N}{M}\right)$
and the desired result comes from Markov’s inequality. ∎
Appendix A A Some LDP results for FK-Ising model at criticality
In this section, we fix $T=T_{c}$ and omit $p_{c}$ from the notation.
The following theorem will be the starting point of this section, which extends the large deviation bound of the largest cluster ([28]) for the Bernoulli percolation model to the FK-Ising model.
Theorem A.1.
There exists a constant $c_{1}>0$ such that
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\geq N^{\frac{15}{8}}x\right)\leq c_{1}\exp\left(-c_{1}^{-1}x^{16}\right)$
holds for all positive integers $N$.
As we remarked above, a continuous version of Theorem A.1 has been shown in [11], and their proof cannot be applied to the discrete settings. So we follow the framework of [28] to prove Theorem A.1.
First, for any positive integers $m<n$, we define $\mathcal{A}(m,n)$ to be the annulus $\Lambda_{n}\setminus\Lambda_{m},$ and define $C(m,n)$ to be the event that there is an open path around $\mathcal{A}(m,n),$ that is, this open path separates $\partial_{ext}\Lambda_{m}$ and $\partial_{int}\Lambda_{n}.$ Also we define $D(m,n)$ to be the event that there exists an open path crossing $\mathcal{A}(m,n),$ or equivalently speaking, connecting $\partial_{int}\Lambda_{n}$ and $\partial_{ext}\Lambda_{m}.$ Finally, define
$\pi(m,n)=\phi^{\mathrm{w},0}_{\mathcal{A}(m,n)}(D(m,n)).$
We point out that here the wired boundary condition on $\mathcal{A}(m,n)$ means that $\partial_{ext}\Lambda_{m}$ and $\partial_{int}\Lambda_{n}$ are wired into one point respectively, but they are not wired together. And we write $\pi(n)=\pi(1,n)$ for short.
Given this notation, we want to check Assumptions 1.1 and 1.2 in [28]. That is
  1. (i)
    There exists a constant $C_{1}>0$ such that for any $0<k<l<m$,
    $\pi(k,l)\pi(l,m)\leq C_{1}\pi(k,m).$
  2. (ii)
    There exists positive constants $C_{2},\alpha>0$ such that for all $n>m\geq 1,$
    $\frac{\pi(n)}{\pi(m)}\geq C_{2}\left(\frac{n}{m}\right)^{-\alpha}.$
Assumption (ii) is for true for $\alpha=\frac{1}{8},$ since in [35] (see also [14, 21, 25]) they have shown that there exists a constant $c>0$ such that for any $n>0$
$c^{-1}n^{-\frac{1}{8}}\leq\pi(n)\leq cn^{-\frac{1}{8}}.$ (A.1)
Then we check assumption (i).
First, if $m\geq 2l\geq 4k,$ then we get that
$\pi(k,m)=\phi^{\mathrm{w},0}_{\mathcal{A}(k,m)}\left(D(k,m)\right)\geq\phi^{\mathrm{w},0}_{\mathcal{A}(k,m)}\left(D(k,m)\mid C(l,2l)\right)\times\phi^{\mathrm{w},0}_{\mathcal{A}(k,m)}\left(C(l,2l)\right).$ (A.2)
Figure 4: An illustration of how $\gamma$ connects the two crossing.
Given the event $C(l,2l)$ happens, there is an open path $\gamma$ in $\mathcal{A}(l,2l)$ separating its inner and outer boundaries, thus $\gamma$ also separates $\partial_{ext}\Lambda_{k}$ and $\partial_{int}\Lambda_{m}.$ As a result, if both $D(k,2l)$ and $D(l,m)$ happen, this two crossing should be connected together by $\gamma$ (See Figure 4 for illustration). Combined with CBC, we have
$\phi^{\mathrm{w},0}_{\mathcal{A}(k,m)}(D(k,m)\mid C(l,2l))\geq\pi(k,2l)\pi(l,m).$ (A.3)
Similarly, we have
$\pi(k,2l)\geq\pi(k,l)\pi\left(\Big{\lfloor}\frac{l}{2}\Big{\rfloor},2l\right)\phi_{\mathcal{A}(k,2l)}^{\mathrm{w},0}\left(C\left(\Big{\lfloor}\frac{l}{2}\Big{\rfloor},2l\right)\right).$ (A.4)
Finally, by the RSW theory (see [21, Theorem 1.1]), we know there is a positive constant $c^{\prime}>0$ such that
$\phi_{\mathcal{A}(m,n)}^{\mathrm{w},0}\left(C(m,n)\right)$ $\geq c^{\prime}\text{ if }n\geq 2m;$
$\phi_{\mathcal{A}(m,n)}^{\mathrm{w},0}\left(D(m,n)\right)$ $\geq c^{\prime}\text{ if }n\leq 8m.$ (A.5)
Now putting (A.2), (A.3), (A.4) and (A) together, we have
$\pi(k,m)\geq(c^{\prime})^{3}\pi(k,l)\pi(l,m).$
Furthermore, (A.4) actually proves that $\pi(k,2l)\geq(c^{\prime})^{2}\pi(k,l)$. Combining with the fact that $\pi(a,b)$ increases with $a$ and decreases with $b$, we can solve the case $m<2l$ and $l<2k$ in a similar but simpler way.
Define
$\mathcal{V}_{n}:=\left\{v\in\Lambda_{n}\,:\,v\leftrightarrow\partial_{int}\Lambda_{2n}\right\}$
to be the set of vertices in $\Lambda_{n}$ which are connected to $\partial_{int}\Lambda_{2n}$. Then [28, Theorem 1.5] and (A.1) show
Lemma A.2.
There is a constant $c_{2}$ such that for all $n>0$
$\left\langle\binom{\left|\mathcal{V}_{n}\right|}{k}\right\rangle_{\Lambda_{2n}}^{\mathrm{w}}\leq(\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}})^{k}.$ (A.6)
Remark A.3.
Although [28] only considers the percolation model, their proof on Theorem 1.5 (in [28]) can be naturally extended to the FK-Ising model.
Corollary A.4.
For $t>n^{-\frac{15}{8}}$, we have
$\left\langle(1+t)^{|\mathcal{V}_{n}|}\right\rangle_{\Lambda_{2n}}^{\mathrm{w}}\leq c_{3}\exp(c_{2}t^{\frac{16}{15}}n^{2}).$
Proof.
We expand $(1+t)^{|\mathcal{V}_{n}|}$ and divide it into two parts at $k=t^{\frac{16}{15}}n^{2}$
[…]
$=\sum_{k=0}^{t^{\frac{16}{15}}n^{2}}\left(t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\right)^{k}+\sum_{k=t^{\frac{16}{15}}n^{2}+1}^{\infty}\left(t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\right)^{k}.$ (A.7)
We now compute the two terms separately. For $k\leq t^{\frac{16}{15}}n^{2}$, we have $t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\leq\frac{c_{2}t^{\frac{16}{15}}n^{2}}{k}$, thus
$\sum_{k=0}^{t^{\frac{16}{15}}n^{2}}\left(t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\right)^{k}\leq\sum_{k=0}^{t^{\frac{16}{15}}n^{2}}\left(\frac{c_{2}t^{\frac{16}{15}}n^{2}}{k}\right)^{k}\leq\sum_{k=0}^{t^{\frac{16}{15}}n^{2}}\frac{(c_{2}t^{\frac{16}{15}}n^{2})^{k}}{k!}\leq\exp(c_{2}t^{\frac{16}{15}}n^{2}).$ (A.8)
For the other term, we group each consecutive $t^{\frac{16}{15}}n^{2}$ elements and apply a union bound since $\left(t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\right)^{k}$ is decreasing, then we obtain
$\sum_{k=t^{\frac{16}{15}}n^{2}+1}^{\infty}\left(t\cdot\frac{c_{2}n^{\frac{15}{8}}}{k^{\frac{15}{16}}}\right)^{k}\leq t^{\frac{16}{15}}n^{2}\sum_{l=1}^{\infty}(l/c_{3})^{-\frac{15}{16}t^{\frac{16}{15}}n^{2}l}\leq t^{\frac{16}{15}}n^{2}\cdot C_{1}.$ (A.9)
Combining (A.8) and (A.9) into (A.7) gives
$\left\langle(1+t)^{|\mathcal{V}_{n}|}\right\rangle_{\Lambda_{2n}}^{\mathrm{w}}\leq C_{2}\exp\left(c_{2}t^{\frac{16}{15}}n^{2}\right).\qed$
Proof of Theorem A.1.
Without loss of generality we only consider the case that $N^{\frac{15}{8}}x\leq N^{2},$ i.e. $x\leq N^{\frac{1}{8}}$ since the volume of any cluster inside $\Lambda_{N}$ is no more than $N^{2}$ Furthermore, we assume $x>\max\{K_{1},K_{2}\}$ where $K_{1},K_{2}>1$ are two large constants depending only on the constants $c_{2}$ and $c_{3}$ in Corollary A.4 to be determined later.
We divide the proof into two parts.
PART 1: Controlling the bulk cluster
In the first part, we show that all the clusters except for the boundary cluster cannot be too large, the proof is really similar to the proof of [5, Proposition 6.3]. In order to be self-contained, we give a short proof here.
Without loss of generality, we may assume that $N=2^{n}$ and define $\mathscr{B}_{l}:=\{\Lambda_{2^{l-1}}(u):u\in(2^{l}\mathbb{Z})^{2}\}$ to be a box partition of $\Lambda_{N}$ with side length $2^{l}.$ Then we write $\mathscr{B}:=\bigcup_{l=1}^{n}\mathscr{B}_{l}.$ Now for a box $\Lambda_{2^{l-1}(u)}\in\mathscr{B}_{l},$ we write $L=2^{l-1}$ for short and define $\mathcal{V}_{L}(u):=\{v\in\Lambda_{2^{l-1}(u)}:v\leftrightarrow\Lambda_{2^{l}(u)}\}.$ We will call the box $\Lambda_{2^{l-1}(u)}$ is crowded if $|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x$ and we have the following claim:
If there exists a cluster in $\Lambda_{N}$ with volume at least $N^{\frac{15}{8}}x$ and it is not the boundary cluster, then at least one box in $\mathscr{B}$ is crowded. The proof is relatively straightforward, use $\mathcal{C}$ to denote this cluster then there must exist an integer $l$ such that $2^{l+1}\leq diam(\mathcal{C})<2^{l+2}$. Then $\mathcal{C}$ will intersect with at most 289 boxes in $\mathscr{B}_{l}$ (this explains why we require that $\mathcal{C}$ is not the boundary cluster). Therefore, at least one of them is crowded since $diam(\mathcal{C})<2^{l+2}$ implies for any $x\in\mathcal{C}$, $x$ connects to some vertex $2^{l}$ away. Therefore, all we need to do is to upper-bound the probability that there is a crowded box in $\mathscr{B}$. Using Corollary A.4 and Markov inequality shows
$\phi_{\Lambda_{2L}(u)}^{\mathrm{w},0}\Big{(}|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x\Big{)}\leq(1+t)^{-\frac{1}{289}N^{\frac{15}{8}}x}\left\langle(1+t)^{|\mathcal{V}_{L}|}\right\rangle^{\mathrm{w}}_{\Lambda_{2L}}$
$\leq(1+t)^{-\frac{1}{289}N^{\frac{15}{8}}x}c_{3}\exp\left(c_{2}t^{\frac{16}{15}}{L}^{2}\right).$ (A.10)
Choosing $t=\min\{\frac{N^{\frac{225}{8}}x^{15}}{K_{1}L^{30}},1\}$ (recall that $1<K_{1}<x$ is the constant defined at the beginning of the proof), then $t>N^{-\frac{15}{8}}$ since $L=2^{l-1}\leq N$. Applying a simple bound $(1+t)^{-\frac{1}{289}N^{\frac{15}{8}}x}\leq\exp(-\frac{t}{2\cdot 289}N^{\frac{15}{8}}x)$ to (A.10), we get that for $\frac{N^{\frac{225}{8}}x^{15}}{K_{1}L^{30}}<1$,
$\phi_{\Lambda_{2L}(u)}^{\mathrm{w},0}\Big{(}|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x\Big{)}\leq c_{3}\exp\left(-\frac{1}{2\cdot 289K_{1}}\frac{N^{30}x^{16}}{L^{30}}+\frac{c_{2}}{K_{1}^{\frac{16}{15}}}\frac{N^{30}}{L^{30}}\right).$
For $\frac{N^{\frac{225}{8}}x^{15}}{K_{1}L^{30}}>1$, we have
$\phi_{\Lambda_{2L}(u)}^{\mathrm{w},0}\Big{(}|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x\Big{)}$ $\leq c_{3}\exp\left(-\frac{1}{2\cdot 289}N^{\frac{15}{8}}x+c_{2}{L}^{2}\right)$
$\leq c_{3}\exp\left(-\frac{1}{2\cdot 289}N^{\frac{15}{8}}x+\frac{c_{2}}{K_{1}^{\frac{1}{15}}}N^{\frac{15}{8}}x\right).$
Let $l_{0}$ denote the maximal integer such that $\frac{N^{\frac{225}{8}}x^{15}}{K_{1}}>2^{30l_{0}}$. Letting $K_{1}$ big enough depending on $c_{2}$ and $c_{3}$, we get that
$\phi_{\Lambda_{2L}(u)}^{\mathrm{w},0}\Big{(}|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x\Big{)}\leq C_{1}\exp\left(-\frac{1}{C_{1}}\frac{N^{30}x^{16}}{L^{30}}\right)~{}~{}\mbox{if }L=2^{l-1},~{}l>l_{0};$ (A.11)
$\phi_{\Lambda_{2L}(u)}^{\mathrm{w},0}\Big{(}|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x\Big{)}\leq C_{1}\exp\left(-\frac{1}{C_{1}}N^{\frac{15}{8}}x\right)~{}~{}\mbox{if }L=2^{l-1},~{}l\leq l_{0}.$ (A.12)
At the same time, there are at most $(\frac{N}{L})^{2}$ elements in $\mathscr{B}_{l}$. The desired result follows from taking a union bound, summing (A.11) and (A.12) over all boxes in $\mathscr{B}$ and recalling $L=2^{l-1}\leq N,$ $x\leq N^{\frac{1}{8}}$:
$\sum_{l=2}^{l_{0}}C_{1}\exp\left(-\frac{1}{C_{1}}N^{\frac{15}{8}}x\right)\cdot\left(\frac{N}{L}\right)^{2}+\sum_{l=l_{0}+1}^{\lfloor\log_{2}N\rfloor}C_{1}\exp\left(-\frac{N^{30}x^{16}}{C_{1}L^{30}}\right)\cdot\left(\frac{N}{L}\right)^{2}\leq C_{2}\exp\left(-C_{2}^{-1}x^{16}\right).$
PART 2: Controlling the boundary cluster
For the second part, we are going to upper-bound the probability that the boundary cluster is too large. We are not able to repeat the proof above since with the wired boundary, the boundary cluster with a small diameter can intersect with a large number of boxes. As a result, we may exploit a multilevel partition of $\Lambda_{N}$ instead of a uniform partition. Without loss of generality, we assume $N=2^{8k}$ and $\lfloor\log_{2}(x)\rfloor=a$. Let $\mathscr{T}_{8k-2}$ denote the set of boxes with side length $2^{8k-1}$ with one corner at the origin $o$. Let $\mathscr{T}_{l}(l=7k+a-1,7k+a,\cdots,8k-3)$ be the collection of boxes with side length $2^{l+1}$ and forms a disjoint contour surrounding boxes in $\mathscr{T}_{l+1}$ (see Figure 5 for illustration).
For convenience, we use $\mathcal{T}_{l}(l=7k+a-1,\cdots,8k-2)$ to denote the union of boxes in $\mathscr{T}_{l}.$ Since the number of points not contained in $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$ is $N^{\frac{15}{8}}2^{a-1}<\frac{N^{\frac{15}{8}}x}{2}$, we only need to consider the intersection of $\mathcal{C}_{*}$ and $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$. Let $C_{3}$ be an absolute constant such that $\sum_{l=7k+a-1}^{8k-2}C_{3}2^{\frac{-8k+l}{2}}\leq 1$. Then we have
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(|\mathcal{C}_{*}|\geq N^{\frac{15}{8}}x\right)\leq\sum_{l=7k+a-1}^{8k-3}\phi_{p_{c},\Lambda_{N}}^{\mathrm{w},0}(|\mathcal{C}_{*}\cap\mathcal{T}_{l}|\geq C_{3}N^{\frac{15}{8}}x\cdot 2^{\frac{-8k+l}{2}}).$ (A.13)
Next we want to control the probability $\phi_{p_{c},\Lambda_{N}}^{\mathrm{w},0}(|\mathcal{C}_{*}\cap\mathcal{T}_{l}|\geq C_{3}N^{\frac{15}{8}}x\cdot 2^{\frac{-8k+l}{2}})$. Here we remark that the factor $2^{\frac{-8k+l}{2}}$ is used to ensure the sequence is summable.
[…]
Figure 5: Multilevel partition of $\Lambda_{N}$
Since $\mathcal{C}_{*}$ is the boundary cluster, we get that for each box $\Lambda_{2^{l}}(u)$ in $\mathscr{T}_{l}$, if it contains a point in $\mathcal{C}_{*}$, then that point is connected to $\Lambda_{2^{l+1}}(u)$. Since the number of boxes in $\mathscr{T}_{l}$ is divisible by $4,$ we can partition the boxes in $\mathscr{T}_{l}$ into $4$ groups $\mathscr{T}_{l}^{1},\cdots\mathscr{T}_{l}^{4}$ (also use $\mathcal{T}_{l}^{1},\cdots,\mathcal{T}_{l}^{4}$ to denote the union of the boxes) such that for any two different boxes $\Lambda_{2^{l}}(u),\Lambda_{2^{l}}(v)$ in a group, we have $\Lambda_{2^{l+1}}(u)\cap\Lambda_{2^{l+1}}(v)=\emptyset$ (see the superscripts in Figure 5 above). Now we can control the number of points connected to the boundary in each group by applying CBC and Corollary A.4 :
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(|\mathcal{C}_{*}\cap\mathcal{T}_{l}^{i}|\geq C_{3}N^{\frac{15}{8}}x\cdot 2^{\frac{-8k+l}{2}}\right)$
[…]
Letting $t=N^{-\frac{15}{8}}x^{15}2^{\frac{15(8k-l)}{16}}/K_{2}<1$ (recall that $1<K_{2}<x$ is the constant in the beginning of the proof and thus, $t>n^{-\frac{15}{8}}$) in (A.14) and noting that $|\mathcal{T}_{l}^{i}|\leq 2^{8k-l}$, we obtain that
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(|\mathcal{C}_{*}\cap\mathcal{T}_{l}^{i}|\geq C_{3}N^{\frac{15}{8}}x\cdot 2^{\frac{-8k+l}{2}}\right)\leq\exp\left(-C_{3}x^{16}2^{\frac{7(8k-l)}{16}}/K_{2}\right)\times c_{3}\exp\left(c_{2}x^{16}/K_{2}^{\frac{16}{15}}\right).$
Letting $K_{2}$ big enough depending on $c_{2}$ and $c_{3}$ gives
$\phi_{\Lambda_{N}}^{\mathrm{w},0}(|\mathcal{C}_{*}\cap\mathcal{T}_{l}^{i}|\geq C_{3}N^{\frac{15}{8}}x\cdot 2^{\frac{-8k+l}{2}})\leq C_{4}\exp(-C_{4}^{-1}x^{16}\cdot 2^{\frac{7(8k-l)}{16}}).$
Combined with (A.13), it completes the proof:
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(|\mathcal{C}_{*}|\geq N^{\frac{15}{8}}x\right)\leq\sum_{l=7k+a-1}^{8k-3}C_{4}\exp\left(-C_{4}^{-1}x^{16}\cdot 2^{\frac{7(8k-l)}{16}}\right)\leq C_{5}\exp\left(-C_{5}^{-1}x^{16}\right).\qed$
For a configuration $\omega\in\Xi$, define $\mathscr{M}(\omega):=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}|\mathcal{C}|^{2},$ to be the summation of the squares of the number of clusters in $\omega.$
Our main goal here is to show the following LDP result:
Theorem A.5.
There exists a positive constant $c_{4}>0$ such that for any $N$ we have
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(\mathscr{M}(\omega)\geq N^{\frac{15}{4}}x\right)\leq c_{4}\exp\left(-c_{4}^{-1}x^{8}\right).$
Proof.
By a simple computation, we have
$\mathscr{M}(\omega)=\sum_{l=1}^{N^{2}}\sum_{\mathcal{C}\in\mathfrak{C},|\mathcal{C}|\geq l}|\mathcal{C}|\leq\sum_{k=0}^{\lfloor 2\log_{2}N\rfloor}2^{k+1}\sum_{\mathcal{C}\in\mathfrak{C},|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|.$
We may divide the summation into three pieces with respect to the level $k$. First for $322^{k}<N^{\frac{7}{4}}x/$ 32, since $\sum_{|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|\leq 4N^{2}$, we get that
$\sum_{2^{k+5}<N^{\frac{7}{4}}}2^{k+1}\sum_{\mathcal{C}\in\mathfrak{C},|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|\leq N^{\frac{15}{4}}x/4.$ (A.15)
Secondly, for $2^{k}\geq N^{\frac{15}{8}}\sqrt{x}$, applying Theorem A.1 gives
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(\sum_{2^{k}\geq N^{\frac{15}{8}}\sqrt{x}}2^{k+1}\sum_{\mathcal{C}\in\mathfrak{C},|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|\geq N^{\frac{15}{4}}x/4\right)$ (A.16)
[…]
The third part is much more complicated. For $N^{\frac{7}{4}}x/32\leq 2^{k}<N^{\frac{15}{8}}\sqrt{x}$, we want to prove that
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(\sum_{\mathcal{C}\in\mathfrak{C},|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|\geq C_{0}N^{\frac{15}{4}}x2^{-k}2^{\frac{k-k_{0}}{2}}\right)\leq C\exp\left(-C^{-1}x^{8}\cdot 2^{k_{0}-k}\right),$ (A.17)
where $k_{0}=\max\{k\in\mathbb{N}:2^{k}<N^{\frac{15}{8}}\sqrt{x}\}$. As in the proof of Theorem A.1, the factor $2^{\frac{k-k_{0}}{2}}$ is introduced to ensure the sequence is summable. We will first consider a dividing procedure of $\Lambda_{N}$. To be precise, we fix an integer $M=M(k)>0$ and consider the following horizontal and vertical line segments inside $\Lambda_{N}$:
$\mathscr{L}_{M}:=\Big{\{}[-N,N]\times\{(2i+1)M\},\{(2i+1)M\}\times[-N,N]:\text{for all }i\text{ with }|(2i+1)M|\leq N\Big{\}}.$
Without loss of generality, we can assume that $(2L+1)M=N$ for some positive integer $L.$ Then $\Lambda_{N}$ is divided into $(2L+1)^{2}$ boxes of size $M\times M$. We will use $\mathscr{B}_{M}$ to denote the set of all these boxes. Furthermore, we define $B(i,j):=[(2i-1)M,(2i+1)M]\times[(2j-1)M,(2j+1)M]$ to denote the box in $\mathscr{B}_{M}.$
Now we consider a new measure, where we artificially restrict all segments in $\mathscr{L}$ to be open. Then the new measure will be a product measure of FK-Ising measure $\phi_{B(i,j)}^{\mathrm{w},0}(\cdot)$ on $\Lambda_{N},$ for each $B(i,j)$, denoted as $\Psi$.
So every cluster $\mathcal{C}$ with $|\mathcal{C}|\geq 2^{k}$ will either be contained in some $B(i,j)$ or intersect with at least one line segment $\mathscr{L}_{M}$. Thus, we can define $\mathscr{C}_{(i,j)}$ to be all the clusters in $B(i,j)$ which are not connected to $\partial_{ext}B(i,j),$ and define $\mathcal{C}^{*}_{(i,j)}$ to be the cluster connected to the boundary in $B(i,j)$. Furthermore, we define
$\mathcal{S}_{(i,j)}(k,M)=\sum_{\mathcal{C}\in\mathscr{C}_{(i,j)},|\mathcal{C}|\geq 2^{k}}|\mathcal{C}|+|\mathcal{C}^{*}_{(i,j)}|.$
By monotonicity, we know $\sum_{\mathcal{C}\geq 2^{k}}|\mathcal{C}|$ is stochastically dominated by
$\mathscr{S}(k,M):=\sum_{B(i,j)\in\mathscr{B}_{M}}\mathcal{S}_{(i,j)}(k,M).$
Now define the event $A_{(i,j)}$ to be the event that the largest cluster in $B(i,j)$ is less than $2^{k},$ by Theorem A.1 we have
$\Psi(A_{(i,j)})\geq 1-c_{1}\exp\left(-\frac{2^{16k}}{c_{1}M^{30}}\right).$
Taking a union bound then gives
$\Psi(A^{*})\geq 1-c_{1}(2L+1)^{2}\exp\left(-\frac{2^{16k}}{c_{1}M^{30}}\right),$
where $A^{*}$ is the intersection of all $A_{(i,j)}.$ Thus, we may choose $M(k)=\lfloor 2^{\frac{8k}{15}-\frac{k_{0}-k}{30}}x^{-\frac{4}{15}}\rfloor$ such that $\frac{2^{16k}}{M^{30}}\geq x^{8}2^{k_{0}-k}$. Note here we restrict $N^{\frac{7}{4}}x/32\leq 2^{k}<N^{\frac{15}{8}}\sqrt{x}$ and thus $M>1$ is well defined. Recall the definition of $k_{0}$, we have $2L+1=\frac{N}{M}\leq C_{1}2^{\frac{17(k_{0}-k)}{30}}$, thus we obtain that
$\Psi(A^{*})\geq 1-C_{2}\exp\left(-C_{2}^{-1}x^{8}2^{k_{0}-k}\right).$
Now it suffices to consider events in $A^{*}$. Given $A_{(i,j)}$ happens, we have that $\mathcal{S}_{(i,j)}(k,M)$ equals to $|\mathcal{C}_{(i,j)}^{*}|$. Furthermore, $A_{(i,j)}$ is a decreasing event, so for any real number $S$, we have by FKG
$\Psi(\{\mathscr{T}(k,M)>S\}\cap A^{*})\leq\Psi(\{\sum|\mathcal{C}^{*}_{(i,j)}|>S\}\cap A^{*}).$
Therefore, it suffices to prove that
$\Psi(\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}N^{\frac{15}{4}}x2^{-k}2^{\frac{k-k_{0}}{2}}\}\cap A^{*})\leq C\exp\left(-C^{-1}x^{8}\right).$
We say a box $B(i,j)$ to be nice if $|\mathcal{C}^{*}_{(i,j)}|\geq V=N^{\frac{7}{4}}x2^{-k}2^{\frac{k-k_{0}}{2}}M^{2}$, denoted as $\mathcal{N}(i,j)$. then applying Theorem A.1 shows
$\phi_{B(i,j)}^{\mathrm{w},0}(\mathcal{N}(i,j))\leq c_{1}\exp(-c_{1}^{-1}[VM^{-\frac{15}{8}}]^{16})=c_{1}\exp(-c_{1}^{-1}V^{16}M^{-30}).$ (A.18)
Recall the definition of $A^{*}$, we get that $|\mathcal{C}^{*}_{(i,j)}|\leq 2^{k}$ for any configuration $\omega\in A^{*}$. Thus, we have
$\Psi\left(\left\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\right\}\cap A^{*}\right)\leq\Phi\left(\text{at least }C_{0}\frac{VN^{2}}{2^{k}M^{2}}~{}\mathcal{N}(i,j)\text{ happens}\right).$ (A.19)
Combining (A.18) and (A.19) shows
$\Psi\left(\left\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\right\}\cap A^{*}\right)$
[…]
Since $M\geq 2^{\frac{8k}{15}-\frac{k_{0}-k}{30}}x^{-\frac{4}{15}}/2$ and $2^{k}\leq 2^{k_{0}}\leq N^{\frac{15}{8}}\sqrt{x}$, we calculate that $V^{16}M^{-30}\geq\frac{N^{2}}{16M^{2}}$. Thus, we obtain that there exists some constant $C_{3}>0$ such that
$\frac{\exp\left(-c_{1}^{-1}V^{16}M^{-30}\right)}{N^{2}/M^{2}}\geq\exp\left(-C_{3}V^{16}M^{-30}\right).$
Thus, we finish the proof of (A.17) by computing the exponent in (A.20) and combining with the fact that $2^{k}\leq 2^{k_{0}}<N^{\frac{15}{8}}\sqrt{x}$. Combining now (A.15), (A.16) and (A.17) gives the desired result. ∎
To finish this subsection, we will give a lower bound for the LDP to show that the exponent $16$ and $8$ in Theorem A.1 and Theorem A.5 are optimal.
Theorem A.6.
There exists a constant $c>0$ such that
$\phi_{\Lambda_{N}}^{\mathrm{w},0}\left(\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\geq N^{\frac{15}{8}}x\right)\geq c\exp\left(-c^{-1}x^{16}\right)$
holds for all positive integers $N$.
The proof of Theorem A.6 is the same as the proof of [28, Theorem 1.4].
Appendix B B Some analytic results
In this section, we present some analytic results on concentration and anticoncentration of random variables.
Lemma B.1.
Let $X$ be a random variable with $\mathsf{var}(X)\geq c_{1}>0$ and ${\mathbb{E}}X^{4}\leq c_{2}$, then there exists a constant $c_{3}>0$ depending only on $c_{1}$ and $c_{2}$ such that there exist $\Omega_{1},\Omega_{2}\subset\Omega$ with $\mathbb{P}(\Omega_{1}),\mathbb{P}(\Omega_{2})\geq c_{3}$ such that
$|X(\omega_{1})-X(\omega_{2})|\geq\frac{\sqrt{c_{1}}}{2},~{}~{}~{}\forall\omega_{1}\in\Omega_{1},\omega_{2}\in\Omega_{2}.$ (B.1)
Proof.
Consider $X_{K}=X\mathbf{1}_{|X|\leq K}$ for any $K>0$. Since ${\mathbb{E}}X^{4}\leq c_{2}$, we compute that for $K>\max\{2c_{1},2\}$
$\mathsf{var}(X_{K})$ $\geq\mathsf{var}(X)-{\mathbb{E}}~{}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K\text{ or }|X^{2}|\geq K\}}$
$=\mathsf{var}(X)-2{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|<K\}}-{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|\geq K\}}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(|X^{1}|+K)^{2}\mathbf{1}_{\{|X^{1}|\geq K\}}-2{\mathbb{E}}[(X^{1})^{2}+(X^{2})^{2}]\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|\geq K\}}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(2|X^{1}|)^{2}\mathbf{1}_{\{|X^{1}|\geq K\}}-4{\mathbb{E}}[(X^{1})^{2}]\mathbf{1}_{\{|X^{1}|\geq K\}}$
$\geq\mathsf{var}(X)-12c_{1}/K^{2},$
where $X^{1},X^{2}$ are sampled independently with the same law as $X$. Then there exists a constant $K_{0}=K_{0}(c_{1},c_{2})$ such that for any $K>K_{0}$, $\mathsf{var}(X_{K})\geq\frac{c_{1}}{2}$.
Now choose $\Omega_{1}$ such that
$\mathbb{P}(\Omega_{1})=\frac{c_{1}}{64K^{2}}\text{ and }\forall\omega\in\Omega_{1},\omega^{\prime}\not\in\Omega_{1},X_{K}(\omega)\geq X_{K}(\omega^{\prime});$
choose $\Omega_{2}$ such that
$\mathbb{P}(\Omega_{2})=\frac{c_{1}}{64K^{2}}\text{ and }\forall\omega\in\Omega_{2},\omega^{\prime}\not\in\Omega_{2},X_{K}(\omega)\leq X_{K}(\omega^{\prime}).$
Therefore we have
$\mathsf{var}(X_{K})={\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}$
$=$ ${\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\{\omega_{1}\in\Omega_{1}\cup\Omega_{2}\text{ or }\omega_{2}\in\Omega_{1}\cup\Omega_{2}\}}+{\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\{\omega_{1},\omega_{2}\in\Omega_{1}^{c}\cap\Omega_{2}^{c}\}}$
$\leq$ $4K^{2}\times\frac{4c_{1}}{64K^{2}}+(\min_{\omega_{1}\in\Omega_{1}}X_{K}(\omega_{1})-\max_{\omega_{2}\in\Omega_{2}}X_{K}(\omega_{2}))^{2}.$
Combined with the fact that $\mathsf{var}(X_{K})\geq\frac{c_{1}}{2}$ ,it yields that $\big{|}\min_{\omega_{1}\in\Omega_{1}}X_{K}(\omega_{1})-\max_{\omega_{2}\in\Omega_{2}}X_{K}(\omega_{2})\big{|}\\ \geq\sqrt{\frac{c_{1}}{4}}.$Thus, we have for any $\omega_{1}\in\Omega_{1},\omega_{2}\in\Omega_{2}$ that
$|X_{K}(\omega_{1})-X_{K}(\omega_{2})|\geq\sqrt{\frac{c_{1}}{4}}.$ (B.2)
Let $K$ large enough such that $c_{2}K^{-4}\leq\frac{c_{1}}{64K^{2}}$ and $\Omega_{1}^{\prime}=\Omega_{1}\cap\{X\leq K\},\Omega_{2}^{\prime}=\Omega_{2}\cap\{X\leq K\}$, then we have
$\min\left\{\mathbb{P}(\Omega_{1}^{\prime}),\mathbb{P}(\Omega_{2}^{\prime})\right\}\geq\min\left\{\mathbb{P}(\Omega_{1}),\mathbb{P}(\Omega_{2})\right\}-c_{2}K^{-4}\geq\frac{c_{1}}{64K^{2}}.$
Since $X=X_{K}$ on $\Omega_{1}^{\prime}\cup\Omega_{2}^{\prime}$, (B.1) follows directly from (B.2). ∎
Remark B.2.
It is easy to see that with ${\mathbb{E}}~{}e^{|X|}\leq c_{2}$ in place of ${\mathbb{E}}X^{4}\leq c_{2}$, the same result still holds.
Corollary B.3.
Let $X$ be a random variable with $\mathsf{var}(X)\geq c_{1}>0$ and ${\mathbb{E}}~{}e^{|X|}\leq c_{2}$,, there exists a constant $c_{3}>0$ depending only on $c_{1},c_{2}$ such that there exist $\Omega^{\prime}\subset\Omega$ with $\mathbb{P}(\Omega^{\prime})\geq c_{3}$ and for any $\omega\in\Omega^{\prime}$, we have
$\big{|}X(\omega)-\ln\left({\mathbb{E}}~{}e^{X}\right)\big{|}\geq\frac{\sqrt{c_{1}}}{4}.$
Lemma B.4.
For any constant $c_{1}>0$, there exists a constant $c_{2}>0$ only relies on $c_{1}$ such that for any random variable $X$ with ${\mathbb{E}}\exp(X)<\infty$ and $\mathsf{var}(X)\geq c_{1}$, we have $\ln\big{(}{\mathbb{E}}\exp(X)\big{)}\geq{\mathbb{E}}X+c_{2}$.
Proof.
Note that $e^{x}\geq 1+x+\frac{x^{2}}{3}$. Thus, we compute
${\mathbb{E}}\exp(X)\geq\exp({\mathbb{E}}X)\left(1+{\mathbb{E}}\left[(X-{\mathbb{E}}X)+\frac{(X-{\mathbb{E}}X)^{2}}{3}\right]\right)=\exp({\mathbb{E}}X)\Big{(}1+\frac{\mathsf{var}(X)}{3}\Big{)}.$
The desired result thus holds since $\mathsf{var}(X)\geq c_{2}$. ∎
Lemma B.5.
Let $X_{1},X_{2},\cdots,X_{n}$ be i.i.d Gaussian variables with mean 00 and variance $1$. Let $Y_{i}=X_{i}^{2}-1$ and $Z_{i}=X_{i}^{4}-6X_{i}^{2}+3$. Then there exists a constant $c>0$ such that for any $a_{i}\geq 0~{}(i=1,2,\cdots,n)$, we have
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}a_{i}Y_{i}\Bigg{|}\geq x\right)\leq c\exp\left(-\frac{x^{\frac{1}{2}}}{c(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{4}}}\right)$ (B.3)
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}a_{i}Z_{i}\Bigg{|}\geq x\right)\leq c\exp\left(-\frac{x^{\frac{1}{4}}}{c(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}\right).$ (B.4)
In particular, if $x\geq 3\sum_{i=1}^{n}a_{i}+\sqrt{\sum_{i=1}^{n}a_{i}^{2}}$, we have
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}a_{i}X_{i}^{4}\Bigg{|}\geq x\right)\leq c\exp\left(-\frac{\left(x-3\sum_{i=1}^{n}a_{i}\right)^{\frac{1}{4}}}{c(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}\right).$ (B.5)
Proof.
The proofs of (B.3) and (B.4) are similar to that of [15, Lemma B.1] (see [31, Lemma 1] for an alternative proof to the first inequality), so we only give a sketch for (B.4) here. We start from controlling the moments of $\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}$. By hypercontractivity for Guassian variables [33] (see also [34, Theorem 1.4.1]), we get that
${\mathbb{E}}\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{p}\leq(p-1)^{2p}\Big{(}{\mathbb{E}}(\sum_{i=1}^{n}a_{i}Z_{i})^{2}\Big{)}^{\frac{p}{2}}=(p-1)^{2p}\Big{(}\sum_{i=1}^{n}a_{i}^{2}\Big{)}^{\frac{p}{2}}.$ (B.6)
By [15, Equation (B.8)], we can control the exponential moment of $\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{1}{4}}$
$\exp(t\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{1}{4}})=\sum_{p=0}^{\infty}{\mathbb{E}}\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{p}{4}}\cdot t^{p}/p!$ $\leq\sum_{p=0}^{\infty}(\sum_{i=1}^{n}a_{i}^{2})^{\frac{p}{8}}(t)^{p}$
$=\frac{1}{1-(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}\cdot t}$
where $t\in(0,\frac{1}{(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}})$. Choosing $t=\frac{1}{2(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}$ and applying an exponential version of the Markov inequality completes the proof of (B.4). The proof of (B.5) comes from the fact that $X_{i}^{4}=3+6Y_{i}+Z_{i}$ and combining with (B.3) and (B.4). ∎
Lemma B.6 (Bernstein’s inequality).
Let $X_{1},X_{2},\cdots,X_{n}$ be independent random variables with a uniform exponential moment bound, i.e., ${\mathbb{E}}\exp(|X_{i}|)\leq C_{1}$ for $1\leq i\leq n$. Then for any $x>0$, there exists a constant $c>0$ relies on $C_{1}$ but not $x$ such that
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}X_{i}-{\mathbb{E}}X_{i}\Bigg{|}\geq x\right)\leq\exp(-\min\{\frac{x^{2}}{nc},\frac{x}{c}\}).$
Proof.
See for example [37, Theorem 2.8.1]. ∎
Appendix C C Some results on FK-Ising model without disorder
In this section, we prove some large deviation results for the FK-Ising model at low-temperature and high-temperature.
Lemma C.1.
Fix $p>p_{c}$. Let $\mathfrak{C}(\omega)$ denote the collection of clusters under the configuration $\omega$ and $\mathcal{C}_{\diamond}$ denote the maximal cluster. Then for any constant $\alpha>0$, there exists a constant $c_{1}=c_{1}(\alpha,p)>0$ such that
$\phi^{\mathrm{w},0}_{p,\Omega}\left(\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{\diamond}\}}|\mathcal{C}|\geq N^{\alpha}\right)\leq c_{1}^{-1}\exp\left(-c_{1}N^{c_{1}}\right)$
holds for any positive integer $N$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$. Furthermore, there exists a constant $c_{2}=c_{2}(p)>0$ such that
$\phi^{\mathrm{w},0}_{p,\Omega}\left(\{\partial_{int}\Omega\subset\mathcal{C}_{\diamond}\}\right)\geq 1-c_{2}^{-1}\exp\left(-c_{2}\sqrt{N}\right)$
holds for any positive integer $N$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$.
Proof.
Let $M=\min\left\{\frac{N}{6},\frac{N^{\frac{\alpha}{2}}}{16}\right\}$. Let $\{B_{i}=\Lambda_{M}(u_{i})\}_{i=1}^{k}$ be a disjoint $2M$-box covering of $\Omega$. Similar to the proof of Lemma 6.10, let $\mathcal{G}(u)$ denote the event that there exists an open circuit in the annuli $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$ and it is connected to $\partial_{int}\Lambda_{M}(u)$ in $\omega.$ Then by the RSW theory, there exists a constant $C_{1}>0$ that only relies on $p$ such that for any boundary condition $\gamma$ on $\partial_{int}\Lambda_{2M}(u_{i}),$
$\phi^{\gamma,0}_{p,\Lambda_{2M}(u_{i})}\left(\mathcal{G}(u_{i})\right)\geq 1-\exp(-C_{1}M).$
Therefore, let $\mathcal{G}=\cap_{i=1}^{k}\mathcal{G}(u_{i})$, then we have by CBC that
$\phi^{\mathrm{w},0}_{p,\Omega}(\mathcal{G})\geq 1-\sum_{i=1}^{k}\phi^{\mathrm{w},0}_{p,\Omega}(\mathcal{G}(u_{i})^{c})\stackrel{{\scriptstyle(*)}}{{\geq}}1-\sum_{i=1}^{k}\phi^{\mathrm{f},0}_{p,\Lambda_{2M}(u_{i})}(\mathcal{G}(u_{i})^{c})$ $\geq 1-\frac{N^{2}}{M^{2}}\exp(-C_{1}M)$
$\geq 1-C_{2}^{-1}\exp(-C_{2}M),$ (C.1)
where the last inequality holds since $M=\min\{\frac{N}{6},\frac{N^{\frac{\alpha}{2}}}{16}\}$. Note that in ()(*)( ∗ ), we do not restrict $\Lambda_{2M}(u_{i})\subset\Omega$.
For $\omega\in\mathcal{G}$, the circuits in the annulus $\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i})$ are connected, and we denote the cluster containing those circuits as $\mathcal{C}_{0}$. In addition, we get that any cluster with diameter larger than $4M$ will be connected to one of the circuits in the annuli $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$, thus $\mathcal{C}_{0}=\mathcal{C}_{\diamond}$ is the maximal cluster in $\omega$ and $\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{\diamond}\}}|\mathcal{C}|<16M^{2}$. In conclusion, we get from (C.1) that
$\phi^{\mathrm{w},0}_{p,\Omega}\left(\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{\diamond}\}}|\mathcal{C}|\geq N^{\alpha}\right)\leq\phi^{\mathrm{w},0}_{p,\Omega}(\mathcal{G}^{c})\leq\exp(-C_{2}M).$
Furthermore, let $\alpha=1$, then for $\omega\in\mathcal{G}$, since $\mathcal{C}_{0}$ is the maximal cluster in $\omega$ and $\partial_{int}\Omega\subset\mathcal{C}_{0}$, we get that $\partial_{int}\Omega\subset\mathcal{C}_{\diamond}$. Combined with (C.1), it yields that
$\phi^{\mathrm{w},0}_{p,\Omega}\left(\{\partial_{int}\Omega\subset\mathcal{C}_{\diamond}\}\right)\geq 1-\exp(-C_{3}\sqrt{N}).\qed$
Lemma C.2.
Fix $p>p_{c}$. Then there exists a constant $c=c(p)>0$ such that for any positive integer $N$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\langle|\mathcal{C}_{*}|\rangle-\langle|\mathcal{C}_{*}\cap\tilde{\mathcal{C}}_{*}|\rangle\geq cN^{2}$
where $\mathcal{C}_{*}$ and $\tilde{\mathcal{C}}_{*}$ are the boundary clusters of two independently sampled configurations under measure $\phi_{p,\Omega}^{\mathrm{w},0}$.
Proof.
We can rewrite the left-hand side as $\langle M\rangle$ where
[…]
Taking expectation w.r.t. $\phi_{p,\Omega}^{\mathrm{w},0}$ we have
$\langle M\rangle=\sum_{x}\phi_{p,\Omega}^{\mathrm{w},0}(x\in\mathcal{C}_{*})\phi_{p,\Omega}^{\mathrm{w},0}(x\not\in\tilde{\mathcal{C}}_{*}).$
Noticing the fact that both probabilities have positive lower bounds only relying on $p$ completes the proof. More precisely, a point has a positive probability of connecting to the boundary since $p>p_{c}$, and a positive probability of not connecting to the boundary due to finite energy property. ∎
Lemma C.3.
Fix $p>p_{c}$. Then there exists a constant $c=c(p)>0$ such that for any positive integer $N$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\phi_{p,\Omega}^{\mathrm{w},0}\Big{(}\big{|}|\mathcal{C}_{\diamond}|-\langle|\mathcal{C}_{\diamond}|\rangle_{p,\Omega}^{\mathrm{w},0}\big{|}\geq N^{1.5}\Big{)}\leq N^{-0.8}.$
Proof.
Let $\mathcal{E}$ denote the event that $\partial_{int}\Omega\subset\mathcal{C}_{\diamond}$ and let $\mathcal{C}_{*}$ denote the boundary cluster. Then we calculate by Lemma C.1 that
$\Big{|}\langle|\mathcal{C}_{\diamond}|-|\mathcal{C}_{*}|\rangle_{p,\Omega}^{\mathrm{w},0}\Big{|}\leq\mathbb{P}(\mathcal{E}^{c})\cdot N^{2}\leq\exp(-C_{1}\sqrt{N})\cdot N^{2}.$
Combined with Lemma C.1 again, it suffices to show that
$\phi_{p,\Omega}^{\mathrm{w},0}\Big{(}\big{|}|\mathcal{C}_{*}|-\langle|\mathcal{C}_{*}|\rangle_{p,\Omega}^{\mathrm{w},0}\big{|}\geq N^{1.5}/2\Big{)}\leq N^{-0.8}/2.$
Next, we calculate the variance of $|\mathcal{C}_{*}|$ under the measure $\phi_{p,\Omega}^{\mathrm{w},0}$.
$\mathsf{var}_{p,\Omega}^{\mathrm{w},0}(|\mathcal{C}_{*}|)$ $=\sum_{x,y\in\Omega}\phi_{p,\Omega}^{\mathrm{w},0}({x,y\in\mathcal{C}_{*}})-\sum_{x,y\in\Omega}\phi_{p,\Omega}^{\mathrm{w},0}({x\in\mathcal{C}_{*}})\cdot\phi_{p,\Omega}^{\mathrm{w},0}({y\in\mathcal{C}_{*}}).$ (C.2)
For any $x,y\in\Omega,$ let $d_{x,y}=\min\{dist(x,y),dist(x,\partial_{int}\Omega),dist(y,\partial_{int}\Omega)\}$ and $M=\max\{d_{x,y}/2,\\ N^{0.1}\}$. Let $\mathcal{E}_{x}$ and $\mathcal{E}_{y}$ denote the event that there is a circuit in $\Lambda_{2M}(x)\setminus\Lambda_{M}(x)$ and $\Lambda_{2M}(y)\setminus\Lambda_{M}(y)$ which is connected to $\partial_{int}\Omega$ respectively. By the proof of Lemma C.1, we get that
$\phi_{p,\Omega}^{\mathrm{w},0}(\mathcal{E}_{x}\cap\mathcal{E}_{y})\geq 1-\exp(-C_{2}M).$ (C.3)
Under $\mathcal{E}_{x}\cap\mathcal{E}_{y}$, the events $x\in\mathcal{C}_{*}$ and $y\in\mathcal{C}_{*}$ are independent, thus we calculate that
$\phi_{p,\Omega}^{\mathrm{w},0}({x,y\in\mathcal{C}_{*}})-\phi_{p,\Omega}^{\mathrm{w},0}({x\in\mathcal{C}_{*}})\cdot\phi_{p,\Omega}^{\mathrm{w},0}({y\in\mathcal{C}_{*}})$
[…]
$-[\phi_{p,\Omega}^{\mathrm{w},0}({x\in\mathcal{C}_{*}}\mid\mathcal{E}_{x})\cdot\left(1-\phi_{p,\Omega}^{\mathrm{w},0}(\mathcal{E}_{x}^{c})\right)]\cdot[\phi_{p,\Omega}^{\mathrm{w},0}({y\in\mathcal{C}_{*}}\mid\mathcal{E}_{y})\cdot\left(1-\phi_{p,\Omega}^{\mathrm{w},0}(\mathcal{E}_{y}^{c})\right)]$
[…]
$+\phi_{p,\Omega}^{\mathrm{w},0}(\mathcal{E}_{x}^{c})+\phi_{p,\Omega}^{\mathrm{w},0}(\mathcal{E}_{y}^{c})$
[…]
Combined with (C.3), it yields that
$\phi_{p,\Omega}^{\mathrm{w},0}({x,y\in\mathcal{C}_{*}})-\phi_{p,\Omega}^{\mathrm{w},0}({x\in\mathcal{C}_{*}})\cdot\phi_{p,\Omega}^{\mathrm{w},0}({y\in\mathcal{C}_{*}})\leq 4\exp(-C_{2}M).$ (C.4)
Plugging (C.4) into (C.2), we get that
$\mathsf{var}_{p,\Omega}^{\mathrm{w},0}(|\mathcal{C}_{*}|)\leq C_{4}N^{2.2}.$ (C.5)
The desired bound thus comes from a Markov inequality. ∎
Lemma C.4.
Fix $p<p_{c}$. Then there exists a constant $c_{1}=c_{1}(p)>0$ such that for any domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and $x>c_{1}$, we have $\phi_{p,\Omega}^{\mathrm{f},0}(\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\geq N^{2}x)\leq c_{1}\exp(-c_{1}^{-1}N\sqrt{x})$ and $\phi_{p,\Omega}^{\mathrm{f},0}(\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\geq N^{2}x)\leq c_{1}\exp(-c_{1}^{-1}\sqrt{N}x^{1/4})$. Furthermore, for any $y\geq N^{0.01}$, we have $\phi_{p,\Omega}^{\mathrm{f},0}(\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\geq y)\leq c_{2}\exp(-c_{2}\sqrt{y})$.
Proof.
We first compute the moments of $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}$. Since $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}=\sum_{u,v\in\Omega}\mathbf{1}_{\{u\longleftrightarrow v\}}$, we compute that for any positive integer $k$
$\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{k}\Big{\rangle}=\Big{\langle}\sum_{u_{i},v_{i}\in\Omega}\prod_{i=1}^{k}\mathbf{1}_{\{u_{i}\longleftrightarrow v_{i}\}}\Big{\rangle}=\sum_{u_{i},v_{i}\in\Omega}\phi_{p,\Omega}^{\mathrm{w},0}\big{(}{\cap_{i=1}^{k}\{u_{i}\longleftrightarrow v_{i}\}}\big{)}.$ (C.6)
Recalling the definition of $\mathtt{Even}_{\{u_{1},\cdots,u_{k},v_{1},\cdots,v_{k}\}}$ from the proof of Lemma 4.3, we get that $\cap_{i=1}^{k}\{u_{i}\longleftrightarrow v_{i}\}\subset\mathtt{Even}_{\{u_{1},\cdots,u_{k},v_{1},\cdots,v_{k}\}}$. By a similar reason to (4.7) and (4.8), we get that
$\sum_{u_{i},v_{i}\in\Omega}\phi_{p,\Omega}^{\mathrm{w},0}\big{(}{\cap_{i=1}^{k}\{u_{i}\longleftrightarrow v_{i}\}}\big{)}\leq C_{1}^{2k}\prod_{j=1}^{k}(N+j)^{2}\leq C_{1}^{2k}(N+k)^{2k}.$ (C.7)
With the moment bounds, we are able to compute an exponential moment for $\sqrt{\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}}$. For any $t>0$, we have
$\Bigg{\langle}\exp\left(t\sqrt{\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}}\right)\Bigg{\rangle}\leq\sum_{k=0}\frac{t^{2k}\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{k}\Big{\rangle}}{(2k)!}+\sum_{k=0}\frac{t^{2k+1}\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{\frac{2k+1}{2}}\Big{\rangle}}{(2k+1)!}.$ (C.8)
By (C.7), we get that
$\frac{t^{2k}\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{k}\Big{\rangle}}{(2k)!}\leq\frac{(2C_{1})^{2k}(N^{2k}+k^{2k})}{(2k)!}\leq\frac{(2C_{1})^{2k}N^{2k}}{(2k)!}+C_{2}^{2k}.$ (C.9)
By Cauchy inequality and (C.7) again, we get that
$\frac{\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{\frac{2k+1}{2}}\Big{\rangle}}{(2k+1)!}\leq\frac{\sqrt{\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{k}\Big{\rangle}\cdot\Big{\langle}\Big{(}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{)}^{k+1}\Big{\rangle}}}{(2k+1)!}$
[…]
Plugging (C.9) and (C.10) into (C.8) and choosing $t=\frac{1}{2C_{2}}$, we get that
$\Bigg{\langle}\exp\left(t\sqrt{\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}}\right)\Bigg{\rangle}$ $\leq\sum_{k=0}\frac{(2C_{1}Nt)^{2k}}{(2k)!}+\frac{(2C_{1}Nt)^{2k+1}}{(2k+1)!}+(C_{2}t)^{2k}+(C_{2}t)^{2k+1}$
$=\exp\left(\frac{C_{1}}{C_{2}}N\right)+2.$ (C.11)
The desired result thus comes from the Markov inequality and letting $c_{1}$ large enough. For $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}$, the result comes in a similar way. For the last criterion, we have by [22] that
$\phi_{p,\Omega}^{\mathrm{f},0}\left(\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\geq y\right)\leq\sum_{\mathsf{dist}(x,y)\geq\sqrt{y}}\phi_{p,\Omega}^{\mathrm{f},0}(x\longleftrightarrow y)\leq N^{2}y\exp(-C_{3}\sqrt{y}).$
The desired result follows from the fact that $y\geq N^{0.01}$. ∎
Corollary C.5.
For any $T>0,$ let $\mathscr{S}=\{\omega:\sum_{\mathcal{C}\in\mathfrak{C}{(\omega)}}|\mathcal{C}|^{2}\leq\epsilon^{-1}N^{3\alpha(T)}\}.$ For any $p>0$, there exists a constant $c>0$ such that
$\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathscr{S})\geq 1-c\exp\left(-c^{-1}\epsilon^{-1}N^{-\alpha(T)}\right).$ (C.12)
Proof.
We divide the proof of this lemma into three parts according to the temperature. When $p=p_{c}$, then the desired result comes from Theorem A.5.
When $p>p_{c}$, then $\alpha(T)=1$. If $\epsilon N\geq 1$, then we have $\epsilon^{-1}N^{-\alpha(T)}\leq 1$. thus we can choose $c$ large enough such that $c\exp(-c)>1$ and then the desired result holds. If $\epsilon N\geq 1$, then we have $\epsilon^{-1}N^{3\alpha(T)}\geq N^{4}$ and thus $\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathscr{S})\geq 1$.
When $p<p_{c}$, then the desired result comes from Lemma C.4. ∎
Lemma C.6.
Fix $p<p_{c}$. Then there exists a constant $c=c(p)>0$ such that for any positive integer $N$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
$\Big{\langle}\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\Big{\rangle}-\Big{\langle}\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\Big{\rangle}\geq cN^{2}$
where $\mathfrak{C}$ and $\tilde{\mathfrak{C}}$ are the collections of all clusters of two independently sampled configurations under measure $\phi_{p,\Omega}^{\mathrm{f},0}$.
Proof.
Using the same idea as the proof of Lemma C.2, we have
[…]
Only considering the pairs of neighboring $x,y$ and applying finite energy property, we get $cN^{2}$ as a lower bound. ∎
Acknowledgements
We thank Jian Ding for several insightful discussions. We thank Xinyi Li and João Maia for reading an early version of this paper and for providing many valuable suggestions. We thank Federico Camia and Jianping Jiang for reminding us the continuous version of Theorem A.1. F. Huang is supported by the Beijing Natural Science Foundation Undergraduate Initiating Research Program (Grant No. QY23006). All authors are partially supported by the National Natural Science Foundation of China (NSFC) through the Tianyuan Mathematics Key Program (Grant No. 12226001) and the NSFC Key Program (Grant No. 12231002).